### v2

# Structure Learning in Gaussian Graphical Models from Glauber Dynamics [^1] This work was supported by the National Science Foundation (NSF) under the award CCF-2048223

Abstract
Gaussian graphical model selection is an important paradigm with numerous applications, including biological network modeling, financial network modeling, and social network analysis. Traditional approaches assume access to independent and identically distributed (i.i.d) samples, which is often impractical in real-world scenarios. In this paper, we address Gaussian graphical model selection under observations from a more realistic dependent stochastic process known as Glauber dynamics. Glauber dynamics, also called the Gibbs sampler, is a Markov chain that sequentially updates the variables of the underlying model based on the statistics of the remaining model. Such models, aside from frequently being employed to generate samples from complex multivariate distributions, naturally arise in various settings, such as opinion consensus in social networks and clearing/stock-price dynamics in financial networks.
In contrast to the extensive body of existing work, we present the first algorithm for Gaussian graphical model selection when data are sampled according to the Glauber dynamics. We provide theoretical guarantees on the computational and statistical complexity of the proposed algorithm’s structure learning performance. Additionally, we provide information-theoretic lower bounds on the statistical complexity and show that our algorithm is nearly minimax optimal for a broad class of problems.
1 Introduction
Graphical models are a powerful class of stochastic models that represent conditional dependence structure between random variables using a graph. They have found extensive applications in fields like bioinformatics [Murad and Brandão, 2021], finance [Zhan et al., 2021], and social network analysis [Farasat et al., 2015], where understanding the dependencies between variables is crucial. An important subclass of graphical models is the Gauss-Markov random field (GMRF) or Gaussian graphical model (GGM), which specifically deals with multivariate Gaussian distributions.
An important challenge in graphical models, including GGMs, is structure learning, where the goal is to identify the graph structure that best fits the observed data. Traditional methods assume access to independent and identically distributed (i.i.d) samples from the underlying distribution (see e.g., [Drton and Maathuis, 2016] for a review of these methods). However, in many practical scenarios, i.i.d samples are often unavailable or hard to obtain and the data is more suitably modeled as stochastic processes that evolve over time [Dennis et al., 2022; Lara et al., 2019]. Our work diverges from traditional methods by addressing these non-i.i.d data settings by focusing on structure learning from data sampled via a natural stochastic process known as Glauber dynamics.
Glauber dynamics (also called the Gibbs Sampler) is an important and well-studied stochastic process that is used to model various types of sequential and dependent data. It is a Markov chain that iteratively updates random variables (nodes) of an underlying undirected graphical model according to their local statistics. Eventually, the distribution of the chain converges to the distribution of the underlying model. It has been used to model the spread of information or diseases across networks [Montanari and Saberi, 2010; Lara et al., 2019], and decisions of agents in coordination games [Auletta et al., 2017] and frequently arises in processes where the stationary distribution is multivariate Gaussian.
In this paper, we study the recovery of the graph structure underlying a Gaussian graphical model given observations or ‘read-outs’ of the Glauber dynamics whose stationary distribution is the multivariate Gaussian. This work extends Bresler et al. [2014], which tackles the case of Ising models. We introduce the first Gaussian graphical model selection algorithm when the data are sampled according to Glauber dynamics, and we provide theoretical guarantees on the computational and statistical complexity of the proposed algorithm’s structure learning performance. Specifically, under certain natural conditions, we show that observing the Glauber dynamics for $\mathcal{O}\left(\frac{d^{2}\operatorname{polylog}p}{\beta_{\rm min}^{4}}\right)$ time is sufficient to estimate a graph structure whose maximum degree is $d$. We also show that for a broad class of problems, we achieve a matching (up to $\operatorname{poly}(d)$ and ${\rm polylog}(p)$ factor) lower bound of $\Omega\left(d^{2}\log\left(p\right)\right)$.
One may consider obtaining $d^{2}\log p$ approximately i.i.d samples by letting Glauber dynamics mix after $\mathcal{O}(p)$ updates and applying traditional structure learning algorithms to learn the graph in $\tilde{\mathcal{O}}(d^{2}p)$ [^2] The notation $g=\tilde{\mathcal{O}}(f)$ implies that there exists constants $c_{1},c_{2}$ such that $g\leq c_{1}f\log(f)^{c_{2}}$ time. However, doing so incurs additional costs (as a function of $p$) due to approximation errors, which need to be carefully accounted for. Additionally, our algorithm is embarrassingly parallelizable – allowing for a computational complexity as low as $\tilde{\mathcal{O}}(d^{2}p)$, since each edge can be estimated in parallel. This is comparable to the best-case (parallel) computational complexity of neighborhood selection style approaches Meinshausen and Bühlmann [2006]. On the other hand, GLASSO Friedman et al. [2008] and PC algorithm Spirtes et al. [2001] have time complexities of $\mathcal{O}(p^{3})$ and $\mathcal{O}(p^{d+2})$, respectively, and are not inherently parallelizable.
The rest of the paper is organized as follows. In Section 2, we introduce notation and preliminaries and formalize the problem statement. In Section 3, we talk about related work, and in Section 4, we provide the intuition behind our algorithm. In Section 5 and Section 6, we present our structure learning algorithm and establish its theoretical results. Finally, in Section 7 and Section 8, we provide a lower bound on the observation time needed for any algorithm to recover the graph structure.
2 Preliminaries & Problem Statement
Notation. Scalar random variables are typically denoted by uppercase letters such as $X$ and $Y$, and the values taken by these random variables are denoted by corresponding lowercase letters such as $x$ and $y$. Random vectors are typically denoted in boldface; e.g., $\mathbf{X}\in\mathbb{R}^{p}$ is a $p-$dimensional random vector. The components of $\mathbf{X}$ and their corresponding sample values are written as $X_{i}$ and $x_{i}$, $i\in[p]$. Furthermore, for an index set $A\subseteq[p]$, $\mathbf{X}_{A}$ denotes the the $|A|-$dimensional subvector of $\mathbf{X}$ that has the variables corresponding indices in $A$ concatenated. Matrices are often denoted with boldface Greek letters in uppercase. For example, $\boldsymbol{\Theta}\in\mathbb{R}^{p\times p}$ is a $p\times p$ matrix. Let $A,B\subseteq[p]$. The submatrix, denoted by $\boldsymbol{\Theta}_{AB}$, is a $|A|\times|B|$ matrix with the elements $\theta_{ab}$ where $a\in A$ and $b\in B$. Typically, continuous-time (and discrete-time stochastic processes) are denoted as a collection $\{Y^{(t)}\}_{t\in\mathbb{R}}$ (and as a sequence $\{X^{(n)}\}_{n\in\mathbb{N}}$) where $t\in\mathbb{R}$ represents time (and $n\in\mathbb{N}$ represents the state number.)
2.1 Gaussian Graphical Models
Let $G=(V,E)$ be a graph on a vertex set $V=[p]=\{1,2,\ldots,p\}$ with an edge set $E\subseteq{V\choose 2}$. For each vertex $i\in[p]$, its neighborhood is defined as $N(i)\triangleq\left\{j:\left\{i,j\right\}\in E\right\}$ and its closure is defined as $\operatorname{cl}(i)\triangleq N(i)\cup\{i\}$. The degree of node $i$ is defined as $d_{i}\triangleq|N(i)|$, while degree of the graph $G$ is $d=\max_{i\in[p]}d_{i}$. A graphical model is an ordered pair $(G,\mathbf{X})$, where $\mathbf{X}$ is a zero-mean $p$-dimensional random vector such that (a) each vertex $i\in[p]$ of the graph is identified with the corresponding component $X_{i}$ of $\mathbf{X}$ and (b) the edge structure of $G$ encodes the conditional (in)dependences of $\mathbf{X}$: $X_{i}\perp\!\!\!\perp X_{j}\mid X_{V\setminus\{i,j\}}\Rightarrow\{i,j\}\notin E$. If this relationship holds, we say that $\mathbf{X}$ is markov with respect to $G$.
$(G,f_{\mathbf{X}})$ is said to be a Gaussian graphical model or a Gauss-Markov random field if $f_{\mathbf{X}}$ is the multivariate normal distribution. That is, there is a symmetric positive definite matrix $\boldsymbol{\Sigma}\in\mathbb{S}^{p}_{++}$ such that
$f_{\mathbf{X}}(\mathbf{x})=\frac{1}{(2\pi)^{p/2}\left\lvert\boldsymbol{\Sigma}\right\rvert^{1/2}}\exp\left(-\frac{1}{2}\mathbf{x}^{T}\boldsymbol{\Sigma}^{-1}\mathbf{x}\right).$
The matrix $\boldsymbol{\Theta}\triangleq\boldsymbol{\Sigma}^{-1}$ is called the precision or concentration matrix of $\mathbf{X}$. An attractive feature of the Gaussian graphical model, which lies at the heart of its broad applicability, is that when $\mathbf{X}$ is Markov with respect to $G$, the following holds333the fact that $\theta_{ij}=0$ implies that the corresponding edge does not exist in the graph requires the so-called faithfulness assumption [Lauritzen, 1996], which is only violated on a set of measure 0; we will assume faithfulness holds here.
$\{i,j\}\notin E\iff(\boldsymbol{\Sigma}^{-1})_{ij}=\boldsymbol{\Theta}_{ij}=0.$ (1)
Another useful feature of the multivariate Gaussian is that it is closed under conditioning [Rue and Held, 2005]. That is, suppose $A\subset V$ and denote by $B$ the set $V\setminus A$ so that $\mathbf{X}=\begin{pmatrix}\mathbf{X}_{A}\\ \mathbf{X}_{B}\end{pmatrix}$ is partitioned by the subvectors $\mathbf{X}_{A}$ and $\mathbf{X}_{B}$ with zero mean and its precision matrix partitioned accordingly,
$\boldsymbol{\Theta}=\begin{pmatrix}\boldsymbol{\Theta}_{AA}&\boldsymbol{\Theta}_{AB}\\ \boldsymbol{\Theta}_{BA}&\boldsymbol{\Theta}_{BB}\end{pmatrix}.$ (2)
Then the distribution of $\mathbf{X_{A}}$ conditioned on $\mathbf{X_{B}}=\mathbf{x}_{B}$ is also Gaussian with mean $\boldsymbol{\mu}_{A|B}$ and precision matrix $\boldsymbol{\Theta}_{A|B}$, where
$\boldsymbol{\mu}_{A|B}=-\boldsymbol{\Theta}^{-1}_{AA}\boldsymbol{\Theta}_{AB}\left(\mathbf{x}_{B}\right)\quad\text{ and }\quad\boldsymbol{\Theta}_{A|B}=\boldsymbol{\Theta}_{AA}.$
Using Schur complements [Horn and Johnson, 2012], we get $\boldsymbol{\Theta}_{AA}^{-1}=\boldsymbol{\Sigma}_{AA}-\boldsymbol{\Sigma}_{AB}\boldsymbol{\Sigma}_{BB}^{-1}\boldsymbol{\Sigma}_{BA}$. From this one can check that, for any $i\in V$, the conditional distribution of $X_{i}$ given $\mathbf{X}_{V\setminus\{i\}}$ is a normal distribution with mean and variance given by
$\mu_{X_{i}|\boldsymbol{X}_{V\setminus\{i\}}}$ $=\sum_{j\in V\setminus\{i\}}-\frac{\theta_{ij}}{\theta_{ii}}X_{j}\text{ and }$
$\sigma^{2}_{X_{i}|\mathbf{X}_{V\setminus\{i\}}}$ $=\sigma_{i}^{2}-\boldsymbol{\Sigma}_{i,V\setminus\{i\}}\boldsymbol{\Sigma}_{V\setminus\{i\},V\setminus\{i\}}^{-1}\boldsymbol{\Sigma}_{V\setminus\{i\},i}.$ (3)
From the above and the relationship between the precision matrix and the graph structure (1), it can be checked that $X_{i}$, given $\mathbf{X}_{V\setminus\{i\}}$, satisfies
$X_{i}=\sum_{j\in N(i)}\beta_{ij}X_{j}+\epsilon_{i},$ (4)
where $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2}_{X_{i}|V\setminus\{i\}})$, $\epsilon_{i}\perp\!\!\!\perp X_{j}$ for $j\in V\setminus\{i\}$, and $\beta_{ij}=-\frac{\theta_{ij}}{\theta_{ii}}$. Indeed, this fact played an important role in Meinshausen and Bühlmann [2006], which is considered a landmark paper on Gaussian graphical model selection.
The above property (4) allows one to define a natural Markov chain on the GGM, called the Glauber dynamics, where the transitions depend only on local conditions.
2.2 Glauber Dynamics & The Data Model
Glauber dynamics, originally introduced in statistical physics by Glauber [1963], has found important applications across various disciplines to study and model the behavior of dynamical systems. Given that it is also equivalent to single-site Gibbs sampling, it is often used as a Markov chain Monte Carlo (MCMC) algorithm to obtain samples from stationary distributions. For a comprehensive overview, the reader may consult [Levin et al., 2006; Martinelli, 1999]. In this section, we provide a self-contained definition of Glauber dynamics as a stochastic process whose stationary distribution is $f_{\mathbf{X}}$, which in turn is associated with a Gaussian graphical model (GGM) $(G,\mathbf{X})$. We will then study the problem of learning the edge structure of $G$, given observations from this stochastic process.
Continuous-time Glauber dynamics is a random process $\{\mathbf{Y}^{(t)}\}_{t\geq\mathbb{R}}$ of $p-$dimensional random vectors $\mathbf{Y}^{(t)}$ defined as follows. The process is initialized with $\mathbf{Y}^{(0)}$, an arbitrary (possibly random) vector. It is then updated at an increasing sequence of random time points $\{S_{n}\}_{n\in\mathbb{N}}\subset\mathbb{R}_{+}$, with $S_{0}=0$, where $n$ represents the state or the round. These update times are distributed such that the time between updates $H_{n}\triangleq S_{n+1}-S_{n}$, called holding times, are independent and exponentially distributed with parameter $p$. Notice that this implies that in one unit of time, there are $p$ updates on average [^4] It can also be checked that this is equivalent to associating a rate $1$ Poisson clock with each random variable, and updating the corresponding random variable at the arrival time according to the appropriate update rule..
During the holding time, the process does not change, and it can be represented by a random vector called the state vector. Therefore, the $n^{th}$ state vector is $\mathbf{X}^{(n)}\triangleq\mathbf{Y}^{(t)}$, $S_{n}\leq t<S_{n+1}$. The updates or transitions between states obey the following rule: At time $S_{n}$, an index $I^{(n)}\in[p]$ is selected uniformly at random. If node $i$ was chosen at round $n$, i.e, $I^{(n)}=i$, the random variable $X_{i}^{(n)}$ is then sampled from $f\left(\mathbf{X}_{i}|\mathbf{X}_{V\setminus\{i\}}=\mathbf{X}^{(n-1)}_{V\setminus\{i\}}\right)$. The values of all other components of the random vector remain unchanged. In the case of multivariate Gaussians the update takes the following form:
$X_{i}^{(n)}$ $=\sum_{j\in N(i)}\beta_{ij}X_{j}^{(n-1)}+\epsilon_{i}^{(n)}\text{and }$ (5)
$X_{k}^{(n)}$ $=X_{k}^{(n-1)},k\neq i,$ (6)
where again $\epsilon_{i}^{(n)}\sim\mathcal{N}\left(0,\sigma^{2}_{X_{i}|\mathbf{X}_{N(i)}}\right)$ is independent of $X_{k}^{(n^{\prime})}$ for all $k\in[p],n^{\prime}\in\mathbb{N}$ except $X_{i}^{(n)}$ and $\beta_{ij}=-\frac{\theta_{ij}}{\theta_{ii}}$. This process is known to converge to corresponding multivariate Gaussian distribution $f_{X}$. We refer the interested reader to Amit [1991], e.g., for convergence results.
As discussed in Section 1, we suppose that this stochastic process generates the data we observe. In particular, we suppose that one observes the continuous-time Glauber dynamics $\{\mathbf{Y}^{(t)}\}$ for $T$ units of time, generating $N=Tp$ updates (in expectation) of the underlying state. In the sequel we will treat the number of updates $N$ as a deterministic quantity given that it concentrates very rapidly to $Tp$. For convenience, we suppose that the dataset includes the time, the identity of the updated node, and the corresponding updated state vector, forming the collection $\mathcal{D}=\left\{\left(\mathbf{X}^{(n)},S_{n},I^{(n)}\right)\right\}_{n=0}^{N}$, where we set $I^{(0)}=1$. The sequence $\{\mathbf{Y}^{(t)}\}_{t=0}^{T}$ is equivalent to the recorded dataset $\mathcal{D}$ (with probability 1) and we will use these interchangeably.
2.3 The Structure Learning Problem
Our goal is to learn the edge set $E$ of the graph $G=(V,E)$ underlying the Gaussian graphical model $\left(G,f_{\mathbf{X}}\right)$, given access to observations $\mathcal{D}=\left\{\left(\mathbf{X}^{(n)},S_{n},I^{(n)}\right)\right\}_{n=0}^{N}\in\left(\mathbb{R}^{p}\times\mathbb{R}\times V\right)^{N}$ from the corresponding Glauber dynamics. We define a structure learning algorithm as a map $\phi:\left(\mathbb{R}^{p}\times\mathbb{R}\times V\right)^{Tp}\rightarrow\binom{V}{2}$ that accepts a dataset and produces an estimate of the graph’s edge structure. The goal is to understand how long the Glauber dynamics need to be observed to confidently recover the graph as a function of various parameters of the problem. Toward this, for any $\delta\in(0,1)$, we define $T_{\ast}(\phi,\delta)$ as the smallest observation time with which the algorithm $\phi$ can be guaranteed to recover the graph structure with probability at least $1-\delta$. That is,
$T_{\ast}(\phi,\delta)$ $=\inf\{T>0:\mathbb{P}\left[\phi(\mathcal{D})=E\right]\geq 1-\delta\}.$ (7)
In Section 5, we introduce an algorithm $\phi$ for which Theorem 1 gives an upper bound on $T_{\ast}(\phi,\delta)$. In Section 7, we provide lower bounds on $\min_{\phi}T_{\ast}(\phi,\delta)$ for a natural and broad family of problems, thereby showing that the algorithm we propose is nearly minimax optimal for this class of problems.
3 Related Work
3.1 Graphical Model Selection from i.i.d Data
In the graphical models literature, much of the work focuses on learning graphical models from i.i.d data. Chow and Liu [1968] consider the learning problem from i.i.d samples on tree graphs. Extensions to this idea on mixed categorical and continuous data in Edwards et al. [2010], to polytrees Chow and Wagner [1973], all assume access to i.i.d samples. Another line of research assuming i.i.d samples specific to learning in Gaussian models is, optimizing $l_{1}$-penalized joint Gaussian log-likelihood Banerjee et al. [2007]; Yuan and Lin [2007] while Meinshausen and Bühlmann [2006] tackled high-dimensional problems through a connection to lasso regression. While this list is necessarily incomplete, we invite the reader to consult Drton and Maathuis [2016]; Zhou [2011] for a more comprehensive treatment of the literature.
3.2 Graphical Model Selection from Dynamics
Recently, there has been growing interest in learning graphical models from dynamical data generated by local Markov chains, particularly Glauber dynamics. The work of Bresler et al. [2014] pioneered this area, focusing on learning Ising models from Glauber dynamics. Dutt et al. [2021] further improved upon this by enhancing node-wise regression algorithms, such as RISE, which were originally designed for the i.i.d. settings. On the other hand, Gaitonde and Mossel [2023] generalized this problem to a broader class of local Markov chains, while Gaitonde et al. [2024] extended the analysis to MRFs with higher-order interactions, beyond binary pairwise models. All these works focus on binary or, at most finite-sized alphabets for the random variables in their models. In contrast, we tackle the problem of structure learning from the well-known dynamical process of Glauber dynamics with Gaussian random variables that take values in $\mathbb{R}$. The fact that these random variables are unbounded introduces unique challenges that make structure learning difficult. We discuss these challenges in detail in the following section and provide intuition for how our algorithm addresses these challenges.
4 Intuition behind our Algorithm
To estimate the edge set of the graph underlying the Gaussian graphical model given observations from Glauber dynamics, our strategy is to construct a hypothesis test for the existence of each potential edge $\{i,j\}\in\binom{V}{2}$. In particular, we will create an estimate of the strength of $\beta_{ij}$, and declare that an edge exists in $G$ if this estimate is above an appropriate threshold. We will next provide some intuition on how we construct this test.
Consider any two nodes, $\{i,j\}\in{V\choose 2}$. Imagine there is a sequence of updates where these two nodes take turns being updated. First, node $i$ gets updated, then node $j$, and then node $i$ again as shown in Fig 1. During this sequence of updates, none of the neighbors of node $i$ (other than possibly node $j$) are updated. More formally, assume there exist integers $0\leq n_{0}\leq n_{1}<n_{2}<n_{3}\leq n_{4}$ such that: (a) $I^{(n)}\notin N(i)\setminus\{j\}$ for all $n$ in the range $n_{0}\leq n\leq n_{4}$, (b) $I^{(n)}=i$ for $n=n_{1},n_{3}$, and $I^{(n)}=j$ for $n=n_{2}$.
[…]
Figure 1: Visualization of the idealized update sequence for vertices $i$ and $j$. At times $n_{1}$ and $n_{3}$, node $i$ is updated, while node $j$ is updated at time $n_{2}$. No updates occur for any other neighbors of $i$ in the interval between $n_{0}$ and $n_{4}$, except possibly $j$.
Under this idealized scenario, given that no neighbor of node $i$ was updated between $n_{0}$ and $n_{4}$, from (5), we expect the change in the value of $X_{i}$ between $n_{1}$ and $n_{3}$ to be solely due to the update of node $j$ at $n_{2}$ (if $j$ is indeed a neighbor of $i$), aside from zero-mean noise terms. This change must depend on (a) the change in $X_{j}$, and (b) the influence of $j$ on $i$; the latter, of course, is captured by $\beta_{ij}$. The following lemma formalizes this relationship and forms the basis for our edge test.
Lemma 1.
Under the idealized conditions on $n_{0},n_{1},n_{2},n_{3},n_{4}$ described above, for any $\mathbf{\bar{x}}\in\mathbb{R}^{|N(i)\setminus j|}$ and $c>0$ we have,
$\mathbb{E}\left[\frac{X_{i}^{(n_{3})}-X_{i}^{(n_{1})}}{X_{j}^{(n_{2})}-X_{j}^{(n_{1})}}\middle|\mathbf{X}_{N(i)\setminus j}^{(n_{0})}=\mathbf{\bar{x}},\left|X_{j}^{(n_{2})}-X_{j}^{(n_{1})}\right|\geq c\right]=\beta_{ij}.$ (8)
Proof.
We write $\mathbb{E}_{\mathbf{\bar{x}},c}\left[\cdot\right]=\mathbb{E}\left[\cdot\middle|\mathbf{X}_{N(i)\setminus j}=\mathbf{\bar{x}},\left|X_{j}^{(n_{2})}-X_{j}^{(n_{1})}\right|\geq c\right]$. Given that node $i$ was updated only at $n_{1}$ and $n_{3}$ while node $j$ was updated at $n_{2}$, we have the following:
$\mathbb{E}_{\mathbf{\bar{x}},c}\left[\frac{X_{i}^{(n_{3})}-X_{i}^{(n_{1})}}{X_{j}^{(n_{2})}-X_{j}^{(n_{1})}}\right]$ =$\stackrel{{\scriptstyle(a)}}{{=}}\mathbb{E}_{\mathbf{\bar{x}},c}\left[\frac{\sum_{k\in N(i)}\beta_{ik}\left(X_{k}^{(n_{3}-1)}-X_{k}^{(n_{1}-1)}\right)}{{X_{j}^{(n_{2})}-X_{j}^{(n_{1})}}}\right]+\mathbb{E}_{\mathbf{\bar{x}},c}\left[\frac{\epsilon_{i}^{(n_{3})}-\epsilon_{i}^{(n_{1})}}{X_{j}^{(n_{2})}-X_{j}^{(n_{1})}}\right]$
=$\stackrel{{\scriptstyle(b)}}{{=}}\mathbb{E}_{\mathbf{\bar{x}},c}\left[\beta_{ij}\frac{X_{j}^{(n_{3}-1)}-X_{j}^{(n_{1}-1)}}{X_{j}^{(n_{2})}-X_{j}^{(n_{1})}}\right]+\mathbb{E}_{\mathbf{\bar{x}},c}\left[\frac{\epsilon_{i}^{(n_{3})}-\epsilon_{i}^{(n_{1})}}{X_{j}^{(n_{2})}-X_{j}^{(n_{1})}}\right]$
=$\stackrel{{\scriptstyle(c)}}{{=}}\beta_{ij},$
where (a) follows from the Glauber update process defined in (5), (b) follows since only node $j$ is updated while other nodes remain unchanged and the last equality follows from the fact that $X_{j}^{(n_{3}-1)}=X_{j}^{(n_{2})}$ and $X_{j}^{(n_{1}-1)}=X_{j}^{(n_{1})}$ since node $j$ is updated only at $n_{2}$ and has remained unchanged at other states. Finally (c) follows from the fact that $\epsilon_{i}^{(n)}$ are zero mean and independent of $X_{j}^{(n)},\forall n\in\mathbb{N},j\neq i$, and the fact that the expectation $\mathbb{E}_{\mathbf{\bar{x}},c}\left[\frac{1}{X_{j}^{(n_{2})}-X_{j}^{(n_{1})}}\right]$ exists (finite) due to conditioning on the event $\left|X_{j}^{(n_{2})}-X_{j}^{(n_{1})}\right|\geq c$ (as $1/X$ is integrable near 00 if $|X|\geq c$). ∎
Note that the above lemma holds whether $\{i,j\}\in E$ or not. This suggests a strategy to estimate the strength of $\beta_{ij}$: find multiple update intervals that satisfy the above condition, and compute the average of the ratio of the changes in between $X_{i}$ and $X_{j}$ values. This is the basis for the construction of our test statistic (see Defintion 4).
However, this naive strategy has some limitations. First, while we can tell whether $i$ and $j$ alternate in updates, there is no way to guarantee that the neighbors of $i$ remain unaffected during the interval in question (since we do not know the graph structure!). To address this, we show how one can choose the length of the update interval appropriately to make sure that the required sequence of updates for nodes $i$ and $j$ is observed frequently enough, while making sure the neighbors of node $i$ are not updated too often; see Lemmas 2-3.
The idea behind our test and this fix is inspired by Bresler et al. [2014]. There are however several challenges specific to the Gaussian case that do not occur in Bresler et al. [2014] for the Ising model case. First, the form of the test statistic is rather different and demands a different analysis technique. That is, the proposed test statistic considers not merely the probability of sign flips, but instead the ratio of the change in the values of samples. Second, this test statistic could be uninformative (even in expectation) if the updates are too large; this is not a concern in the Ising model case, where all random variables are bounded. Finally, the variance of this test statistic can be unbounded, potentially overwhelming the signal if the changes in $X_{j}$ are too small. Note that without a uniform lower bound on the absolute value of denominator of the test statistic, its expectation does not exist. Our algorithm and analysis carefully deal with these while also ensuring that the test statistic retains enough power to recover the underlying graph.
5 The Algorithm
From the discussion above, our strategy is to construct a test statistic to detect edges in $G$ by estimating the conditional expectation in Lemma 1. We know this quantity is only useful when the conditions described in Figure 1 occur. Towards constructing an estimator, we will suppose that we choose a small $\tau>0$, and divide the length $T$ of the Glauber dynamics process into $k_{max}\triangleq\lfloor\frac{T}{\tau}\rfloor$ intervals of length $\tau$. We will next show that with a careful choice of $\tau$, we can enforce that our desired events occur frequently enough to get a usable test statistic.
First, for any pair of vertices $\{i,j\}\in[p]$ and $k\in[k_{\rm max}]$, we define the event $A_{ij}^{k}$ which records whether the updates of the pair $\{i,j\}$ occur in the required order in the $k$-th time interval. We write $I_{t}$ to represent the identity of the last update node on or before time $t$.
Definition 1 (Event $A_{ij}^{k}$).
For $k\in[k_{max}]$, let the interval $[(k-1)\tau,k\tau)$ be divided into thirds as below
$W_{1}=[(k-1)\tau,(k-2/3)\tau),$
$W_{2}=[(k-2/3)\tau,(k-1/3)\tau),\text{ and }W_{3}=[(k-1/3)\tau,k\tau).$
For $i,j\in[p]$, we define
[…]
$\bigcap\bigg{\{}\exists t\in W_{3}:I_{t}=i,\forall t\in W_{3}:I_{t}\neq j\bigg{\}}.$
Whenever this event occurs, we update our test statistic with the ratio of changes as suggested by (8). We next show that the event occurs with a non-vanishing probability, and therefore we know that its impact on the test statistic is controllable by the choice of $\tau$.
Lemma 2 (Probability of event $A_{ij}^{k}$).
For $k\in[k_{max}]$, $i,j\in[p]$ we have
$\mathbb{P}\left[A_{ij}^{k}\right]=\left[(1-e^{-\tau/3})e^{-\tau/3}\right]^{3}\triangleq q.$
Next, we define event $D_{ij}^{k}$ for a node $i$ and time interval $k$ which records whether the neighbors of node $i$ are updated in the interval $[(k-1)\tau,k\tau)$.
Definition 2 (Event $D_{ij}^{k}$).
For any $k\in[k_{max}]$ and $\{i,j\}$ pair we define
$D_{ij}^{k}=\left\{I_{t}\notin N(i)\setminus\{j\},\forall(k-1)\tau\leq t<k\tau\right\}.$
As we noted earlier, we cannot observe whether this event occurs in the data or not. However, in the following lemma, we show that we can control how often this event occurs by controlling the parameter $\tau$. The smaller $\tau$ is, the more the chances that $D_{ij}^{k}$ occurs. This stands in tension with the requirement that $A_{ij}^{k}$ occurs frequently for every $j\in[p]$. As part of our analysis, we will identify a value for $\tau$ that balances these requirements.
Lemma 3.
The probability of event $D_{ij}^{k}$ for any $k\in[k_{max}]$ we have $\mathbb{P}\left[D_{ij}^{k}\right]\geq e^{-\tau d}.$
Following the discussion in Lemma 1, our test statistic will be computed as the ratio of changes in the values of coordinates when the appropriate $A_{ij}^{k}$ event occurs. There is, however, a catch. Given that the $D_{ij}^{k}$ events are not observed, the expected value of the test statistic could depart from its idealized value (i.e., the edge strength) since we cannot condition on the appropriate $D_{ij}^{k}$. While we can make the probability of $D_{ij}^{k}$ not occurring as small as we want, if the test statistic is unbounded, this can essentially render the test powerless; that is, the test may not be able to detect the presence or absence of edges even in expectation. Moreover, even if the presence of an edge is detectable in expectation, the variance of the test statistic could be very large if the size of the update of a variable at any point is too small. Note that Bresler et al. [2014], Dutt et al. [2021], and Gaitonde et al. [2024] do not have to deal with this as the Ising model is bounded and nicely behaved. To account for these eventualities, we next define two events and establish controls on their probabilities.
Lemma 4 (Event C).
For any $\delta>0$ and sufficiently large $p$, there exist constants $C_{1},C_{1}^{\prime},C_{1}^{\prime\prime}>0$ such that
$\mathbb{P}\left[\max_{i\in[p],t<C_{1}^{\prime}p^{C_{1}^{\prime\prime}}}\left|Y_{i}^{(t)}\right|\leq C_{1}\sigma_{\rm max}\sqrt{\log\left(\frac{p}{\delta}\right)}\right]\geq 1-\delta/2.$
For a fixed $\delta$, we will define $y_{\rm max}\triangleq C_{1}\sigma_{\rm max}\sqrt{\log\left(\frac{p}{\delta}\right)},$ and $C_{\delta}\triangleq\left\{\max_{i\in[p],t<C_{1}^{\prime}p^{C_{1}^{\prime\prime}}}\left|Y_{i}^{(t)}\right|\leq y_{\textrm{max}}\right\},$ and define the appropriate conditional measure and expectation as $\mathbb{P}_{C}[\cdot]=\mathbb{P}\left[\cdot|C_{\delta}\right]$ and $\mathbb{E}_{C}\left[\cdot\right]=\mathbb{E}\left[\cdot|C_{\delta}\right]$. Letting $T_{\rm max}\triangleq C_{1}^{{}^{\prime}}p^{C_{1}^{\prime\prime}}$, the bound above ensures that the observed trajectories of the Glauber dynamics are bounded for all time $t<T_{\rm max}$. As we prove in Theorem 1, we only need to observe the process for $\Omega\left(\frac{d^{2}\operatorname{polylog}p}{\beta_{\rm min}^{4}}\right)$ time, and for appropriately chosen constants, the event $C_{\delta}$ holds for significantly longer. This forms one part of our strategy to control the test statistic in light of the possible unboundedness discussed above.
Next, we also need to ensure that each change in the Glauber dynamics process is significant enough so that the expectation and variance of the test statistic is bounded. Toward this, for $j\in[p],k\in[k_{\rm max}]$, we introduce the following event $B_{j}^{k}$. This event occurs when $Y_{j}$ changes by at least $\sigma_{\rm min}\triangleq\min_{i\in[p]}\sigma_{i|V\setminus{\{i\}}}$. We assume that we know (a lower bound on) $\sigma_{\rm min}$.
Definition 3 (Event $B_{j}^{k}$).
Given $\{\mathbf{Y}^{(t)}\}_{t=0}^{T}$, for $k\in[k_{\rm max}]$, $j\in[p]$ and $\sigma_{\rm min}$ as defined above, we define
$B_{j}^{k}=\left\{\left\lvert Y_{j}^{(k-1/3)\tau}-Y_{j}^{(k-1)\tau}\right\rvert\geq\sigma_{\rm min}\right\}.$
We next bound the probability of the $B_{j}^{k}$ events conditioned on the $A_{ij}^{k},C_{\delta},D_{ij}^{k}$ events. The value of the process at an infinitesimal time before some $t$ will be written as $\mathbf{Y}^{(t)-}$.
Lemma 5 (Conditional probability of event $B_{j}^{k}$).
Under the conditional measure $\mathbb{P}_{C}$ with $\delta>0$, $k\in[k_{\rm max}]$, $i,j\in[p]$ and any vector $\mathbf{\bar{y}}\in\mathbb{R}^{p}$ with $|\mathbf{\bar{y}}|\leq y_{\rm max}$, we have
$\mathbb{P}_{C}\left[B_{j}^{k}\middle|A_{ij}^{k},D_{ij}^{k},\mathbf{Y}^{(k-1)\tau-}=\mathbf{\bar{y}}\right]>0.78-\delta/2\triangleq b.$
We can now define our test statistic based on Lemma 1 and the event we have defined and Lemma 6 shows that our test statistic is bounded almost surely.
Definition 4 (Test statistic).
For any $i,j\in[p]$ we define
$T_{ij}=\sum_{k=1}^{k_{\rm max}}\mathds{1}_{\{A_{ij}^{k}\cap B_{j}^{k}\}}\frac{\Delta Y_{i}^{k}}{\Delta Y_{j}^{k}}.$
Here $\Delta Y_{i}^{k}=Y_{i}^{k\tau-}-Y_{i}^{(k-2/3)\tau-}$ and $\Delta Y_{j}^{k}=Y_{j}^{(k-1/3)\tau}-Y_{j}^{(k-1)\tau}$. We will let $T_{ij}^{(k)}$ denote the $k$-th term in the above summation.
The events $C_{\delta}$ and $B_{j}^{k}$ allow us to control the behavior of the test statistic by bounding the test statistic shown by Lemma 6 given below and bounding the variance as shown in Lemma 9, addressing two challenges raised in Section 4.
Lemma 6.
For $k\in[k_{\rm max}]$ and $i,j\in[p]$ the $k$-th term of the test statistic is bounded almost surely with respect to the conditional measure $\mathbb{P}_{C}$. That is,
[…]
We now present our structure learning algorithm. The algorithm takes as parameters the dataset $\mathcal{D}$, the interval length $\tau$, a threshold $\rho$, and a margin or error $\delta$. First, it calculates $k_{\rm max}$, the number of $\tau$-length intervals in the dataset, and checks if the high-probability event $C_{\delta}$ occurs in the dataset. If not, it terminates the procedure and returns an empty edgeset (error). Otherwise, the algorithm iterates over all pairs of nodes $\{i,j\}\in\binom{V}{2}$, and, for each pair, computes the test statistic $T_{ij}$. If this test statistic over $k_{\rm max}$ intervals exceeds the threshold $\rho$, we declare an edge between nodes $i$ and $j$.
Algorithm 1
1:Input: $\mathcal{D}$ (dataset), $\tau$ (interval length), $\rho$ (threshold), $\delta$ (confidence parameter)
2:Let $\hat{E}=\emptyset$ and $k_{\rm max}=\lfloor\frac{T}{\tau}\rfloor$
3:if $\mathds{1}_{C_{\delta}}=1$
[…]
5:    if $\left|\frac{1}{k_{\rm max}}\sum_{k=1}^{k_{\rm max}}T_{ij}^{(k)}\right|\geq\rho$, add $\{i,j\}$ to $\hat{E}$.
6:return $\hat{E}$
6 Characterization of the Performance of Algorithm 1
We state our assumptions and regularity conditions that must be satisfied and then we provide theoretical guarantees on the computational and statistical complexity of the proposed algorithm’s structure learning performance.
We suppose that the edge strengths ($\beta_{ij}$ for some nodes $i,j$) are bounded away from zero and the edges are statistically distinguishable from non-edges.
Assumption A1 (Bounded edge strength).
For all pair of edges $\{i,j\}\in E$, the partial regression coefficients $\beta_{ij}$ are bounded by $\beta_{\rm max}$ and $\beta_{\rm min}$.
Further, we assume that no node has zero or arbitrarily large variance.
Assumption A2 (Bounded variance).
For all $i\in[p]$, the marginal variance is bounded above $\sigma_{\rm max}^{2}$ and the conditional variance of $X_{i}$ given $X_{V\setminus\{i\}}$ is bounded below by $\sigma_{\rm min}^{2}$.
In many real-world processes like opinion dynamics DeGroot [1974] where nodes take a weighted average of their neighbor’s value, large sample values do not cascade uncontrollably and rather attenuate spatiotemporally. We assume this property holds for our case and the next assumption ensures this.
Assumption A3 (Bounded degree and Sample decay).
We assume for a graphs $G$ associated $(G,f_{\mathbf{X}})$ to have bounded degree $d$, such that $d\beta_{\rm max}<1$.
Theorem 1.
Consider observations from Glauber dynamics $\{\mathbf{Y}^{(t)}\}_{t=0}^{T}$ associated with a Gaussian graphical model $(G,f_{\mathbf{X}})$. Under the assumptions  A1-A3, for any $\delta\in(0,1/2)$ and sufficiently large $p$ there exists constants $C_{2},C_{3}$ and $C_{4}$ such that if the observation time $T$ satisfies ($T_{\rm max}$ is as in Lemma 4)
$T_{\rm max}\geq T\geq\frac{C_{2}}{(\beta_{\rm min}\sigma_{\rm min})^{4}}d^{2}\log\left(\frac{p}{\delta}\right)^{3},$ (9)
and the interval length $\tau$ and threshold $\rho$ are set as
$\tau=d^{-1}\left[\frac{C_{3}\sqrt{\log\left(\frac{p}{\delta}\right)}}{\sigma_{\rm min}\beta_{\rm min}}+1\right]^{-1},\qquad\rho=C_{4}\tau d\frac{\sqrt{\log\left(\frac{p}{\delta}\right)}}{\sigma_{\rm min}},$
then Algorithm 1 recovers the true edge set with probability at least $1-\delta$.
Remark 1. Our algorithm is guaranteed to recover the true edge set given $\mathcal{O}(\frac{d^{2}\operatorname{polylog}p}{\beta_{\rm min}^{4}})$ Glauber updates per node. Notice that this quantity is an upper bound on $T_{\ast}$ defined in Section 2. The dependence on $\beta_{\rm min}^{4}$ is possibly improvable with a different algorithmic choice and is an interesting avenue for future work.
Remark 2. The computational complexity of Algorithm 1 is $\mathcal{O}\left(d^{2}p^{3}\operatorname{polylog}p/\beta_{\rm min}^{4}\right)$ as the algorithm iterates over (${p\choose 2}$ pairs of vertices. It is worth noting that in the high-dimensional regime, our algorithm recovers the edge set in a significantly shorter observation time than it takes Glauber dynamics to mix which is $\Omega(p)$.
We provide a brief summary of the proof strategy here and provide a full proof of Theorem 1 in Appendix A.1.
Proof Strategy. To bound the probability of error of Algorithm 1, we first condition on the event that $C_{\delta}$ happens (which we choose to be a high-probability event). Then, for each pair of vertices $\{i,j\}$, we first establish (in Lemma 7) that the test statistic proposed is able to distinguish, in expectation, between the cases where $\{i,j\}$ are connected and where they are not; this requires a careful selection of the interval $\tau$ and allows us to pick an appropriate threshold $\rho$. Finally, we establish the finite data performance of our algorithm using a martingale concentration argument that leverages our analysis involving the events $B_{j}$ and $A_{ij}$. It is worth noting that without carefully accounting for the above events, the underlying test statistics is not be a (sub)martingale and is not well-behaved.
7 Observation Time Lower Bounds
In this section, we present minimax lower bounds on the observation time $T$ required to learn the Gaussian graphical model from the Glauber dynamics $\{\mathbf{Y}^{(t)}\}_{t=0}^{T}$. We start by introducing a natural class of graphical models for which we can provide these lower bounds. We then define the minimax risk and present the information-theoretic lower bound.
7.1 Classes of Graphical Models
Let $\mathcal{G}_{p,d}$ denote the class of undirected graphs with vertex set $[p]$ and maximum degree $d\leq p$. With a slight abuse of notation, for a fixed $\lambda>0$, we will let $\mathcal{G}_{p,d}(\lambda)$ denote a set of positive definite matrices $\Theta$ that satisfy the following properties:
  1. 1.
    For any $\Theta\in\mathcal{G}_{p,d}(\lambda)$, there exists a unique $G=([p],E)\in\mathcal{G}_{p,d}$ for which $\Theta_{E^{c}}=0$. That is, the sparsity pattern of $\Theta$ is captured by a graph in $\mathcal{G}_{p,d}$.
  2. 2.
    For every $\Theta\in\mathcal{G}_{p,d}(\lambda)$, the minimum partial regression coefficient satisfies $\beta_{\min}=\min_{\{i,j\}\in E}\frac{|\Theta_{ij}|}{\Theta_{ii}}\geq\lambda.$
Notice that every $\Theta$ in $\mathcal{G}_{p,d}(\lambda)$ can be associated with a multivariate normal distribution with zero mean and covariance matrix $\Theta^{-1}$. Therefore $\mathcal{G}_{p,d}(\lambda)$ is a collection of Gaussian graphical models whose underlying graphs have bounded degrees and bounded normalized edge strengths. This is the collection we work with.
7.2 Graph Decoders and Maximum Probability of Error
Let $\mathcal{X}^{(n)}=(\mathbf{X}^{(1)},\dots,\mathbf{X}^{(n)})\subset\mathbb{R}^{p}$ represent $n$ samples from a distribution from the class of distributions $\mathcal{G}_{p,d}(\lambda)$. At each round $i\in[n]$, the sample $\mathbf{X}^{(i)}$ evolves according to the Glauber dynamics defined in (5). Let $\mathcal{I}^{(n)}=(I^{(1)},\dots,I^{(n)})$ be the sequence of node update indices chosen uniformly at random, where only one component $X^{(k)}_{I^{(k)}}$ is updated at each round. The goal is to estimate the underlying graph $G\in\mathcal{G}_{p,d}$ from the observed sequence $\{\mathcal{X}^{(n)},\mathcal{I}^{(n)}\}$. Note that dataset $\mathcal{D}$ defined in Section 2.3 also included the update times $\{S_{n}\}$. However, the Poisson process associated with the update times, is independent of the identity of the updated node $I^{(k)}$ and the value of the update $X_{I^{(k)}}^{(k)}$. Due to this, the analysis of lower bounds can equivalently be done without $S_{n}$. For details, we refer the reader to the remark in the proof of Lemma 10. Therefore in the lower bound analysis we consider the dataset as the set containing pairs of only $\mathbf{X}^{(n)}$ and $I^{(n)}$.
To prove our minimax lower bound, we will consider graph decoders of the form $\Psi:\{\mathcal{X}^{(n)},\mathcal{I}^{(n)}\}\rightarrow\mathcal{G}_{p,d}$ which take an input data sequence and produce an estimate $\widehat{G}\in\mathcal{G}_{p,d}$. The performance is evaluated using the 0-1 loss function $\mathds{1}[\widehat{G}\neq G]$, indicating whether the estimated graph differs from the true graph.
Let $\mathcal{P}=[\mathcal{G}_{p,d}(\lambda)\times\text{Unif}[p]]^{n}$ be the underlying class of product distributions on the input data sequence. The minimax risk is defined as:
$\mathcal{M}_{n,p,d}(\mathcal{G}_{p,d}(\lambda))$ $=\min_{\Psi}\max_{P^{(n)}\in\mathcal{P}}\mathbb{E}_{D^{(n)}}\left[\mathds{1}[\Psi\neq G]\right]$
$=\min_{\Psi}\max_{\Theta(G)\in\mathcal{G}_{p,d}(\lambda)}\mathbb{P}_{\Theta(G)}[\Psi\neq G].$
Here, $P^{(n)}$ represents the $n$-fold product distribution and the probability measure is taken over the data distribution $\Theta(G)$ for graphs $G\in\mathcal{G}_{p,d}$. The minimax risk represents the minimum error achieved by the best algorithm $\widehat{G}$ in the worst-case scenario, where distributions on graphs are hard to distinguish. Our main goal is to understand the scaling laws of the triplet $(n,p,d)$, where $n$ is the number of updates or samples observed, $d$ is the maximum degree of the graph, and $p$ is the number of vertices such that the minimax risk $\mathcal{M}_{n,p,d}$ vanishes asymptotically.
An alternative viewpoint, which we incorporate to maintain consistency with the achievability theorem’s style, is to consider the time interval $T$ over which the continuous-time Glauber process $\{\mathbf{Y}^{(t)}\}_{t=0}^{T}$ was observed. These two perspectives are equivalent in the sense that $T=\frac{n}{p}$, as there are, on average, $p$ updates over a unit interval of time. Recall that in Section 2.2, we treat the number of samples observed over $T$ time units to be deterministic. Therefore we denote it with lower case $n$. Moreover, as noted earlier, $\{\mathbf{Y}^{(t)}\}$ and the input data sequence $\{\mathcal{X}^{(n)},\mathcal{I}^{(n)}\}$ can be used interchangeably.
7.3 Minimax Lower Bound
We now state the lower bound on the observation time $T$ for graphical model selection.
Theorem 2.
For the class $\mathcal{G}_{p,d}(\lambda)$, if the observation time $T$ for the continuous-time Glauber dynamics $\{\mathbf{Y}^{(t)}\}_{t=0}^{T}$ satisfies
$T\geq\frac{\log\left(\frac{pd}{4}\right)\delta\left[\frac{1-2(1-\delta)}{2(1-\delta)}\right]}{K\lambda^{2}\log p},\quad\text{with}\ \delta\in(0,1/2),$ (10)
for some universal constant $K$, then for any estimator $\widehat{G}$, the maximum probability of error satisfies
$\max_{\Theta(G)\in\mathcal{G}_{p,d}(\lambda)}\mathbb{P}_{\Theta(G)}[\widehat{G}\neq G]\leq\frac{1}{2}.$ (11)
We refer the reader to Appendix A.2 for a complete proof of the above theorem. Let $\lambda=\beta_{\rm min}$ where $\beta_{\rm min}$ is the minimum edge strength as defined in Assumption A1. As a consequence of Theorem 2, if the minimum edge strength $\beta_{\rm min}$ decays at the rate $\mathcal{O}\left(\frac{1}{d\sqrt{\log p}}\right)$, then the lower bound on the observation time becomes $T=\Omega(d^{2}\log p)$. In this regime, the sample complexity $T$ from Theorem 1 matches the information-theoretic lower bound up to a factor of $\operatorname{poly}(d)$ and ${\rm polylog}(p)$. Hence, the proposed structure learning algorithm is nearly optimal.
8 Discussion
We present the first algorithm for learning the structure of Gaussian graphical models when the data are sampled according to a stochastic process called Glauber dynamics, which is a natural model for dependent systems like signal dynamics on graphs. We establish bounds on the observation time required for our algorithm to reconstruct the graph structure, which scales in problem parameters as $\Omega\left(\frac{d^{2}\log^{3}p}{\beta_{\rm min}^{4}}\right)$, and show that it is nearly minimax optimal in a range of problems, that is, matching upto ${\rm poly}(d)$ and ${\rm polylog}(p)$ factors. Although it looks like the sample complexity of our algorithm matches the bound one would obtain by applying “standard” model selection methods to mixed Glauber dynamics samples, this requires careful analysis and likely results in more factors that depend on $p$ to account for the approximation errors. Our algorithm is also embarrassingly parallelizable and offers computational efficiency in this non-i.i.d setting. Further, we expect that the dependence on certain parameters (e.g., $\beta_{\rm min}$, ${\rm polylog}(p)$) can be improved by leveraging alternative algorithmic or analysis techniques. This is a promising avenue for future work.
Appendix
Appendix A Proofs of Theorems
In this section, we will state and prove Theorem 1 and Theorem 2 under the assumptions A1-A3 specified in Section 6. The statements and proofs of all the lemmas used to establish the theorems are provided in Appendix B.
A.1 Theorem 1: Statement & proof
The first theorem establishes achievability results for the observation time needed for model selection by characterizing the performance of Algorithm 1.
Theorem 1.
Consider observations from Glauber dynamics $\{\mathbf{Y}^{(t)}\}_{t=0}^{T}$ associated with a Gaussian graphical model $(G,f_{\mathbf{X}})$. Under the assumptions A1-A3, for any $\delta\in(0,1/2)$ and sufficiently large $p$ there exists constants $C_{2},C_{3}$ and $C_{4}$ such that if the observation time $T$ satisfies ($T_{\rm max}$ is as in Lemma 4)
$T_{\rm max}\geq T\geq\frac{C_{2}}{(\beta_{\rm min}\sigma_{\rm min})^{4}}d^{2}\log\left(\frac{p}{\delta}\right)^{3},$ (12)
and the interval length $\tau$ and threshold $\rho$ are set as
$\tau=d^{-1}\left[\frac{C_{3}\sqrt{\log\left(\frac{p}{\delta}\right)}}{\sigma_{\rm min}\beta_{\rm min}}+1\right]^{-1},\qquad\rho=C_{4}\tau d\frac{\sqrt{\log\left(\frac{p}{\delta}\right)}}{\sigma_{\rm min}},$
then Algorithm 1 recovers the true edge set with probability at least $1-\delta$.
Proof.
To bound the probability of error, which is the probability of the event that the estimated edge set is not equal to the true edge set, we first condition on the high probability event $C_{\delta}$ and its complement as defined in Section 5. As the algorithm returns an empty edge set when $C_{\delta}^{c}$ occurs, the error probability, in this case, is $\mathbb{P}\left[C_{\delta}^{c}\right]<\delta/2$. In the sequel, we analyze the probability of error when the event $C_{\delta}$ has occurred and show that it is also less than $\delta/2$.
Considering each pair $\{i,j\}$ separately, we consider the cases where $\{i,j\}\in E$ and $\{i,j\}\notin E$. We show in Lemma 7 that for an appropriately chosen $\tau$, when the edge does not exist, the expectation of the test statistic is bounded around zero and is bounded away from zero when the edge exists. Furthermore, the bounds are such that there is a separation between them, allowing us to set a threshold between the cases.
Error occurs, when the algorithm claims the existence of an edge between nodes $i,j$ when there is none and vice versa. However, due to separation, with enough samples, we can guarantee the recovery of the true edge set. We use a Bernstein-style martingale concentration inequality, Lemma 8 to establish bounds on the probability of errors. To do this, we need our test statistic to be bounded, almost surely, and we need a bound on the conditional variance of the submartingale. Conditioning on the event $C_{\delta}$ and using events $B_{ij}$ allows us to do both. We expect to be able to strengthen these results using a sub-Gaussian version of martingale concentration [Shamir, 2011] and using an appropriate stopping time construction, however, we leave this for future work.
We begin implementing the above strategy below. We will use the same notation $\mathbb{P}_{C}[\cdot]=\mathbb{P}\left[\cdot|C_{\delta}\right]$ as in Section 5 for the conditional measure with respect to the event $C_{\delta}$. Fix any $\delta>0$. Then the probability of error decomposes as follows
$\mathbb{P}\left[\hat{E}\neq E\right]$ $=\mathbb{P}_{C}\left[\hat{E}\neq E\right]\mathbb{P}\left[C_{\delta}\right]+\mathbb{P}_{C^{c}}\left[\hat{E}\neq E\right]\mathbb{P}\left[C_{\delta}^{c}\right]$
$\leq\mathbb{P}_{C}\left[\hat{E}\neq E\right]+\mathbb{P}\left[C_{\delta}^{c}\right].$ (13)
We show in Lemma 4 that the event $C_{\delta}$ is a high probability event and therefore $\mathbb{P}\left[C_{\delta}^{c}\right]<\delta/2$ holds. Now we focus on bounding the other term in (A.1) which is the probability of error conditioned on $C_{\delta}$.
Using union bound we write the error as a sum of the probability of error for every $\{i,j\}\in\binom{V}{2}$, that is
$\mathbb{P}_{C}\left[\hat{E}\neq E\right]$ $=\mathbb{P}_{C}\left[\exists\{i,j\}\in\binom{V}{2}:\mathds{1}_{\hat{E}}(i,j)\neq\mathds{1}_{E}(i,j)\right]$
$\leq\sum_{\{i,j\}}\mathbb{P}_{C}\left[\mathds{1}_{\hat{E}}(i,j)\neq\mathds{1}_{E}(i,j)\right]$
=$=\sum_{\{i,j\}}\mathds{1}_{E}(i,j)\mathbb{P}_{C}\left[\mathds{1}_{\hat{E}}(i,j)=0\right]+\mathds{1}_{E^{c}}(i,j)\mathbb{P}_{C}\left[\mathds{1}_{\hat{E}}(i,j)=1\right]$
=$=\sum_{\{i,j\}}\Bigg{(}\mathds{1}_{E}(i,j)\mathbb{P}_{C}\left[\left\lvert\frac{\sum_{k=1}^{k_{\rm max}}T_{ij}^{k}}{k_{\rm max}}\right\rvert\leq\rho\right]+\mathds{1}_{E^{c}}(i,j)\cdot\mathbb{P}_{C}\left[\left\lvert\frac{\sum_{k=1}^{k_{\rm max}}T_{ij}^{k}}{k_{\rm max}}\right\rvert>\rho\right]\Bigg{)}.$ (14)
The last equality follows from the definition of Algorithm 1 as it asserts the existence of an edge $\{i,j\}$ if and only if $\frac{\sum_{k=1}^{k_{\rm max}}T_{ij}^{k}}{k_{\rm max}}>\rho$. Here $\rho$ is a threshold parameter taken as input by the algorithm. Next, we show how to choose this threshold parameter.
Recall that from Section 5 under the event $C_{\delta}$ we have, $\max_{i\in[p],t\leq T_{max}}\left\lvert Y_{i}^{(t)}\right\rvert\leq y_{\rm max}$. As discussed in the proof strategy, selecting a threshold involves selecting an appropriate interval length $\tau$ such that a separation exists. In Lemma 7 we establish that if we take
$\tau=d^{-1}\left[\frac{6y_{\rm max}}{b\sigma_{\rm min}\beta_{\rm min}}+1\right]^{-1},$
then we have the following for every $k\in[k_{max}]$ and $i,j\in[p]$:
  1. 1.
    if $\{i,j\}\notin E$, then $\left\lvert\mathbb{E}_{C}\left[T_{ij}^{k}\middle|\mathcal{F}_{k-1}\right]\right\rvert\leq\eta$ where $\eta=\frac{2y_{\rm max}}{\sigma_{\rm min}}q\tau d>0$, and
  2. 2.
    if $\{i,j\}\in E$, then $\operatorname{sign}(\beta_{ij})\mathbb{E}_{C}\left[T_{ij}^{k}\middle|\mathcal{F}_{k-1}\right]\geq\operatorname{sign}(\beta_{ij})\eta^{\prime}$ where $\eta^{\prime}=\frac{4y_{\rm max}}{\sigma_{\rm min}}q\tau d>0$
where we define the sigma algebra $\mathcal{F}_{k-1}=\sigma\left(\left\{\mathbf{Y}^{(t)}\right\}_{0\leq t<(k-1)\tau}\right)$. We therefore take the threshold to be
$\rho=\frac{3\eta}{2}=\frac{3y_{\rm max}}{\sigma_{\rm min}}q\tau d.$
We will now consider the cases of the pair $\{i,j\}$ being connected and not connected separately and bound the terms in (A.1).
Case 1: $\{i,j\}\in E$. We begin by observing that the sequence $\{Z_{k}\}_{k\geq 1}$ defined below is adapted to the filtration $\{\mathcal{F}_{k}\}_{k\geq 1}=\left(\sigma(\{\mathbf{Y}^{(t)}\}_{t=0}^{k\tau})\right)_{k\geq 1}$ as $Z_{k}$ is $\mathcal{F}_{k}$-measureable.
$\{Z_{k}\}_{k\geq 1}=\left\{\sum_{l=1}^{k}\operatorname{sign}(\beta_{ij})T_{ij}^{k}-k\eta^{\prime}\right\}_{k\geq 1}$
Specifically, $T_{ij}^{k}$ is deterministic given the data in the $k$-th interval. Furthermore, from our choice of $\tau$, the following holds due to Lemma 7:
$\mathbb{E}_{C}\left[Z_{k}\middle|\mathcal{F}_{k-1}\right]$ $=\mathbb{E}_{C}\left[Z_{k-1}+\operatorname{sign}(\beta_{ij})T_{ij}^{k}-\eta^{\prime}\middle|\mathcal{F}_{k-1}\right]$
$\geq Z_{k-1}$
as $\mathbb{E}_{C}\left[\operatorname{sign}(\beta_{ij})T_{ij}^{k}-\eta^{\prime}\middle|\mathcal{F}_{k-1}\right]>0$. Therefore, $\{Z_{k}\}$ is the submartingale adapted to the filtration $\{\mathcal{F}_{k}\}_{k\geq 1}$. We will next show that given enough samples we can control the probability of error. To do this we note that the sequence has bounded difference (a.s.) which follows from Lemma 6. Indeed,
$\left\lvert\operatorname{sign}(\beta_{ij})T_{ij}^{k}-\eta^{\prime}\right\rvert\leq\left\lvert T_{ij}^{k}\right\rvert+\eta^{\prime}\leq\frac{2y_{\rm max}}{\sigma_{\rm min}}(1+2q\tau d)=c.$
Next, we show in Lemma 9 we have a bound on the variance of the appropriately conditioned submartingale sequence, i.e $\operatorname{var}\left(Z_{k}|\mathcal{F}_{k-1}\right)\leq\frac{4y_{\rm max}^{2}q}{\sigma_{\rm min}^{2}}=s^{2}$. The probability of error can now be written as
$\mathbb{P}_{C}\left[\left\lvert\frac{\sum_{k=1}^{k_{\rm max}}T_{ij}^{k}}{k_{\rm max}}\right\rvert\leq\rho\right]$ $\leq\mathbb{P}_{C}\left[\frac{\sum_{k=1}^{k_{\rm max}}\operatorname{sign}(\beta_{ij})T_{ij}^{k}}{k_{\rm max}}\leq\rho\right].$
Multiplying $k_{\rm max}$, then subtracting $\eta^{\prime}k_{\rm max}$ both sides allows us to apply the Bernstein style martingale concentration inequality (stated in Lemma 8) with $Z_{0}=0$. This gives us the following.
$\mathbb{P}_{C}\left[Z_{k}\leq(\rho-\eta^{\prime})k_{\rm max}\right]$ $=\mathbb{P}_{C}\left[Z_{k}\leq-\frac{\eta k_{\rm max}}{2}\right]$ (15)
$\leq\exp\left(-\frac{\eta^{2}k_{\rm max}^{2}/4}{2k_{\rm max}s^{2}+c\eta k_{\rm max}/6}\right)$ (16)
$=\exp\left(-\frac{k_{\rm max}}{8(s/\eta)^{2}+2/3(c/\eta)}\right).$ (17)
Taking the union bound over $p^{2}$ edges, setting $\delta/2$ as the upper bound and solving for $k_{\rm max}$we get,
$\left[8\left(\frac{s}{\eta}\right)^{2}+\frac{2}{3}\left(\frac{c}{\eta}\right)\right]\log\left(\frac{2p^{2}}{\delta}\right)$ $\leq 16\left(\frac{1}{q\tau^{2}d^{2}}+\frac{1}{q\tau d}\right)\log\left(\frac{p}{\delta}\right)$
$\leq 32\left(\frac{1}{q\tau^{2}d^{2}}\right)\log\left(\frac{p}{\delta}\right)$
$\leq k_{\rm max}.$
The first inequality follows from the fact that $\log\left(\frac{2p^{2}}{\delta}\right)<2\log\left(\frac{p}{\delta}\right)$ when $\delta<1/2$ and using $(1):\left(\frac{c}{\eta}\right)=1/q\tau d+2<2/q\tau d$ as both $q<1,\tau d<1$ and $(2):\left(\frac{s}{\eta}\right)^{2}=\frac{q}{(q\tau d)^{2}}=\frac{1}{q\tau^{2}d^{2}}$. To convert this bound for $k_{\rm max}$ to $T$ we use $k_{\rm max}=\lfloor\frac{T}{\tau}\rfloor$ we get
$32\left(\frac{1}{q\tau d^{2}}\right)\log\left(\frac{p}{\delta}\right)$ $\stackrel{{\scriptstyle(a)}}{{\leq}}864e\left(\frac{1}{\tau^{4}d^{2}}\right)\log\left(\frac{p}{\delta}\right)$
$\stackrel{{\scriptstyle(b)}}{{\leq}}864e\cdot d^{2}\left(\frac{3y_{\rm max}}{b\beta_{\rm min}\sigma_{\rm min}}+1\right)^{4}\log\left(\frac{p}{\delta}\right)$
$\stackrel{{\scriptstyle(c)}}{{\leq}}\frac{864e6^{4}}{b^{4}}\cdot d^{2}\left(\frac{y_{\rm max}}{\beta_{\rm min}\sigma_{\rm min}}\right)^{4}\log\left(\frac{p}{\delta}\right)$
$\stackrel{{\scriptstyle(d)}}{{\leq}}\frac{C_{2}}{(\beta_{\rm min}\sigma_{\rm min})^{4}}d^{2}\log\left(\frac{p}{\delta}\right)^{3}<T.$
Recall that $q=[(1-e^{-\tau/3})e^{-\tau/3}]^{3}$. For (a) we use $1-e^{-x}>xe^{-a}$ for $x\in[0,a]$ with $\tau<1/2$. Therefore $q>\tau^{3}e^{-0.5}e^{-\tau}/27$ and $e^{\tau}<e^{1/2}$. For (c), recall the choice for $\tau=d^{-1}\left(\frac{3y_{\rm max}}{b\beta_{\rm min}\sigma_{\rm min}}+1\right)^{-1}<d^{-1}\left(\frac{6y_{\rm max}}{b\beta_{\rm min}\sigma_{\rm min}}\right)^{-1}$. Finally for (d) using $y_{\rm max}=C_{1}\sigma_{\rm max}\sqrt{\log\left(\frac{p}{\delta}\right)}$ we get the final bound for $T$.
Case 2: $\{i,j\}\notin E$. Consider the sequences (1): $\{Z_{k}\}_{k\geq 1}=\{\sum_{l=1}^{k}T_{ij}^{l}-k\eta\}_{k\geq 1}$ and (2): $\{\tilde{Z}_{k}\}_{k\geq 1}=\{\sum_{l=1}^{k}T_{ij}^{l}+k\eta\}_{k\geq 1}$. From Lemma 7, it again follows that these are supermartingale and submartingale sequences respectively, adapted to the filtration $\{\mathcal{F}_{k}\}_{k\geq 1}$ as defined before. Using $\eta^{\prime}>\eta$, we indeed have the same bound as in case 1 for the difference sequences associated with $\{Z_{k}\}$ and $\{\tilde{Z}_{k}\}$. That is,
$\left\lvert T_{ij}^{k}-\eta\right\rvert\leq\left\lvert T_{ij}^{k}+\eta\right\rvert\leq\left\lvert T_{ij}^{k}\right\rvert+\eta^{\prime}\leq\frac{2y_{\rm max}}{\sigma_{\rm min}}+\eta^{\prime}=c.$
From Lemma 9, we have $\operatorname{var}\left(Z_{k}|\mathcal{F}_{k-1}\right)=\operatorname{var}\left(\tilde{Z_{k}}|\mathcal{F}_{k-1}\right)=\frac{4y_{\rm max}^{2}}{\sigma_{\rm min}^{2}}q=s^{2}$.
The probability of error in this case can be written as,
$\mathbb{P}_{C}\left[\left\lvert\frac{\sum_{k=1}^{k_{\rm max}}T_{ij}^{k}}{k_{\rm max}}\right\rvert>\rho\right]$ $=\mathbb{P}_{C}\left[Z_{k}>(\rho-\eta)k_{\rm max}\right]+\mathbb{P}_{C}\left[\tilde{Z}_{k}<(\eta-\rho)k_{\rm max}\right]$
and from Lemma 8 we have,
$\mathbb{P}_{C}\left[\left\lvert\frac{\sum_{k=1}^{k_{\rm max}}T_{ij}^{k}}{k_{\rm max}}\right\rvert>\rho\right]$ $\leq 2\exp\left(-\frac{\eta^{2}k_{\rm max}^{2}/4}{2k_{\rm max}s^{2}+c\eta k_{\rm max}/6}\right)=2\exp\left(-\frac{k_{\rm max}}{8(s/\eta)^{2}+2/3(c/\eta)}\right).$ (18)
Taking the union bound over $\binom{p}{2}\leq\frac{p^{2}}{2}$ edges, setting $\delta/2$ as the upper bound and solving for $k_{\rm max}$ we get the same bound for T as in case 1.
A.2 Theorem 2: Statement & Proof
In this section, we present the proof for the minimax lower bounds on the observation time $T$ required to learn the Gaussian graphical model from the Glauber dynamics $\{\mathbf{Y}^{(t)}\}_{t=0}^{T}$. We start by introducing a natural class of graphical models for which we can provide these lower bounds. We then define the minimax risk and state Theorem 2, followed by a detailed presentation of its proof.
Let $\mathcal{G}_{p,d}$ denote the class of undirected graphs with vertex set $[p]$ and maximum degree $d\leq p$. With a slight abuse of notation, for a fixed $\lambda>0$, we will let $\mathcal{G}_{p,d}(\lambda)$ denote a set of positive definite matrices $\Theta$ that satisfy the following properties:
  1. 1.
    For any $\Theta\in\mathcal{G}_{p,d}(\lambda)$, there exists a unique $G=([p],E)\in\mathcal{G}_{p,d}$ for which $\Theta_{E^{c}}=0$. That is, the sparsity pattern of $\Theta$ is captured by a graph in $\mathcal{G}_{p,d}$.
  2. 2.
    For every $\Theta\in\mathcal{G}_{p,d}(\lambda)$, the minimum partial regression coefficient satisfies $\beta_{\min}=\min_{\{i,j\}\in E}\frac{|\Theta_{ij}|}{\Theta_{ii}}\geq\lambda.$
Notice that every $\Theta$ in $\mathcal{G}_{p,d}(\lambda)$ can be associated with a multivariate normal distribution with zero mean and covariance matrix $\Theta^{-1}$. $\mathcal{G}_{p,d}(\lambda)$ is, therefore, a collection of Gaussian graphical models whose underlying graphs have bounded degrees and bounded normalized edge strengths. This is the collection we work with.
A.2.1 Graph Decoders and Maximum Probability of Error
Let $\mathcal{X}^{(n)}=(\mathbf{X}^{(1)},\dots,\mathbf{X}^{(n)})\subset\mathbb{R}^{p}$ represent $n$ samples from the class of distributions $\mathcal{G}_{p,d}(\lambda)$. At each round $i\in[n]$, the sample $\mathbf{X}^{(i)}$ evolves according to the Glauber dynamics defined in (5). Let $\mathcal{I}^{(n)}=(I^{(1)},\dots,I^{(n)})$ be the sequence of node update indices chosen uniformly at random, where only one component $X^{(k)}_{I^{(k)}}$ is updated at each round. The goal is to estimate the underlying graph $G\in\mathcal{G}_{p,d}$ from the observed sequence $\{\mathcal{X}^{(n)},\mathcal{I}^{(n)}\}$. Note that dataset $\mathcal{D}$ defined in Section 2.3 also included the update times $\{S_{n}\}$. However, the Poisson process associated with the update times, is independent of the identity of the updated node $I^{(k)}$ and the value of the update $X_{I^{(k)}}^{(k)}$. Due to this, the analysis of lower bounds can equivalently be done without $S_{n}$. For details, we refer the reader to the remark in the proof of Lemma 10. Therefore in the lower bound analysis, we consider the dataset as the set containing pairs of only $\mathbf{X}^{(n)}$ and $I^{(n)}$.
To prove our minimax lower bound, we will consider graph decoders of the form $\Psi:\{\mathcal{X}^{(n)},\mathcal{I}^{(n)}\}\rightarrow\mathcal{G}_{p,d}$ which take an input data sequence and produce an estimate $\widehat{G}\in\mathcal{G}_{p,d}$. The performance is evaluated using the 0-1 loss function $\mathds{1}[\widehat{G}\neq G]$, indicating whether the estimated graph differs from the true graph.
Let $\mathcal{P}=[\mathcal{G}_{p,d}(\lambda)\times\text{Unif}[p]]^{n}$ be the underlying class of product distributions on the input data sequence. The minimax risk is defined as:
$\mathcal{M}_{n,p,d}(\mathcal{G}_{p,d}(\lambda))$ $=\min_{\Psi}\max_{P^{(n)}\in\mathcal{P}}\mathbb{E}_{D^{(n)}}\left[\mathds{1}[\Psi\neq G]\right]$
$=\min_{\Psi}\max_{\Theta(G)\in\mathcal{G}_{p,d}(\lambda)}\mathbb{P}_{\Theta(G)}[\Psi\neq G].$ (19)
Here, $P^{(n)}$ represents the $n$-fold product distribution and the probability measure is taken over the data distribution $\Theta(G)$ for graphs $G\in\mathcal{G}_{p,d}$. The minimax risk represents the minimum error achieved by the best algorithm $\widehat{G}$ in the worst-case scenario, where distributions on graphs are hard to distinguish. Our main goal is to understand the scaling laws of the triplet $(n,p,d)$—where $n$ is the number of updates or samples observed, $d$ is the maximum degree of the graph, and $p$ is the number of vertices—such that the minimax risk $\mathcal{M}_{n,p,d}$ vanishes asymptotically.
An alternative viewpoint, which we incorporate to maintain consistency with the achievability theorem’s style, is to consider the time interval $T$ over which the continuous-time Glauber process $\{\mathbf{Y}^{(t)}\}_{t=0}^{T}$ was observed. These two perspectives are equivalent in the sense that $T=\frac{n}{p}$, as there are, on average, $p$ updates over a unit interval of time. Recall that in Section 2.2, we treat the number of samples observed over $T$ time units to be deterministic. Therefore we denote it with lower case $n$. Moreover, as noted earlier, $\{\mathbf{Y}^{(t)}\}$ and the input data sequence $\{\mathcal{X}^{(n)},\mathcal{I}^{(n)}\}$ can be used interchangeably.
A.2.2 Fano’s Method
[…]
In order to construct a restricted ensemble, we start with a base graph $G_{0}\in\mathcal{G}_{\text{p,d}}$ consisting of $\left\lfloor\frac{p}{d+1}\right\rfloor$ cliques of size $d+1$. Let $\{C_{1},\ldots,C_{N}\}$ denote the cliques where $N=\left\lfloor\frac{p}{d+1}\right\rfloor$. The inverse covariance matrix $\Theta_{0}$ corresponding to the graph $G_{0}$ is given by
$\Theta_{0}(\lambda)=I+\lambda\sum_{i=1}^{N}\mathds{1}_{C_{i}}\mathds{1}_{C_{i}}^{\top}.$ (20)
where $\lambda$ is the minimum edge strength, $I\in\mathbb{R}^{p\times p}$ is the identity matrix and $\mathds{1}_{C_{i}}$ is the indicator vector corresponding to set $C_{i}$. Let $\mathcal{V}=\{1,2,\ldots,M\}$ denote the index set of the restricted ensemble with $M=|\mathcal{V}|$. For each $i\in V$, the graph $G_{i}$ in the ensemble is generated by removing a single edge from $G_{0}$. The corresponding set of inverse covariance matrices is given by $\{\Theta_{0},\ldots,\Theta_{M}\}$, where for each
$\Theta_{i}=\Theta_{0}-\lambda\left(\mathds{1}_{e_{i}}\mathds{1}_{e_{i}}^{\top}-I_{e_{i}}\right),$
[…]
$p_{\text{err}}(\Psi)$ $=\max_{v\in[M]}\mathbb{P}_{\Theta_{v}(G)}[\Psi(\{\mathcal{X}^{(n)},\mathcal{I}^{(n)}\})\neq G]$ (21)
$\stackrel{{\scriptstyle(a)}}{{\geq}}1-\frac{I(V;\{\mathcal{X}^{(n)},\mathcal{I}^{(n)}\})+1}{\log M}$ (22)
where $(a)$ follows from Fano’s inequality (corollary 2.6 in Tsybakov [2008]). We evaluate the mutual information $I(V;\{\mathcal{X}^{(n)},\mathcal{I}^{(n)}\})$ by upper bounding it by the KL-divergence between the models in the ensemble.
We now state the lower bound on the observation time $T$ for graphical model selection.
Theorem 2.
For the class $\mathcal{G}_{p,d}(\lambda)$, if the observation time $T$ for the continuous-time Glauber dynamics $\{\mathbf{Y}^{(t)}\}_{t=0}^{T}$ satisfies
$T\geq\frac{\log\left(\frac{pd}{4}\right)\delta\left[\frac{2(1-\delta)-1}{2(1-\delta)}\right]}{K\lambda^{2}\log p},\quad\text{with}\ \delta\in(0,1/2),$ (23)
for some universal constant $K$, then for any estimator $\widehat{G}$, the maximum probability of error satisfies
$\max_{\Theta(G)\in\mathcal{G}_{p,d}(\lambda)}\mathbb{P}_{\Theta(G)}[\widehat{G}\neq G]\leq\frac{1}{2}.$ (24)
Note on Parity Change: It is important to note that while the result in the main paper is presented as $\left[-\frac{2(1-\delta)-1}{2(1-\delta)}\right]$, the correct expression is $\left[\frac{2(1-\delta)-1}{2(1-\delta)}\right]$, as shown here. This change in parity does not affect the overall correctness of the result, and the expression in the appendix should be considered the accurate form.
Proof.
We begin by noting that throughout the proof, the Glauber process $\{Y\}_{t=0}^{T}$ and the observation sequence $\{\mathcal{X}^{(n)},\mathcal{I}^{(n)}\}$ are used interchangeably. Additionally, we alternate between the concept of $n$, which represents the total number of Glauber updates within the time interval $[0,T]$, and the total observation time $T$, with the understanding that $T=n/p$. From Fano’s inequality, the maximum probability of error $p_{\text{err}}(\Psi)$ can be lower bounded as,
$p_{\text{err}}(\Psi)$ $\geq 1-\frac{I\left(V;\{\mathcal{X}^{(n)},\mathcal{I}^{(n)}\}\right)+1}{\log M}$ (25)
Let $C_{\delta,n}=\left\{\max_{i\in[n]}|X^{(i)}|<C\sqrt{\frac{\log n}{\delta}}\right\}$, for any $\delta\in(0,1/2)$, denote the event that the maximum value of the Glauber process is bounded by $C\sqrt{\frac{\log n}{\delta}}$, where $C$ is an appropriately chosen constant. We derive the minimax lower bound conditioned on this high-probability event. This conditioning allows us to avoid cases where the Glauber updates may become unbounded, which occur with extremely low probability. Therefore, even with this conditioning, the derived minimax lower bound remains both valid and practical.
Then,
$p_{\text{err}}(\Psi)$ $\geq p(C_{\delta,n})p_{\text{err}}(\Psi|C_{\delta,n})$ (26)
$\stackrel{{\scriptstyle(a)}}{{\geq}}(1-\delta)p_{\text{err}}(\Psi|C_{\delta,n})$ (27)
$\stackrel{{\scriptstyle(b)}}{{\geq}}(1-\delta)\left[1-\frac{I\left(V;\{\mathcal{X}^{(n)},\mathcal{I}^{(n)}\}|\mathds{1}\{C_{\delta,n}\}=1\right)+1}{\log M}\right]$ (28)
where (a) follows since $p(C_{\delta,n})\geq 1-\delta$ (see Lemma 4) and (b) follows from the conditional Fano’s inequality. The goal is to determine the minimum value of $T$ such that $p_{\text{err}}(\Psi)\leq\frac{1}{2}$. It follows from Fano’s inequality that if $p_{\text{err}}(\Psi)\leq\frac{1}{2}$ then $p_{\text{err}}(\Psi|C_{\delta,n})\leq\frac{1}{2(1-\delta)}$. In the remainder of the proof, we determine the minimum value of $T$ for which $p_{\text{err}}(\Psi|C_{\delta,n})\leq\frac{1}{2(1-\delta)}$. Towards this, we begin by computing the conditional mutual information $I(V;\{\mathcal{X}^{(n)},\mathcal{I}^{(n)}\}|\mathds{1}\{C_{\delta,n}\}=1)$. For the ease of notation, we denote the event $\mathds{1}\{C_{\delta,n}\}=1$ by $Z=1$, where $Z=\mathds{1}\{C_{\delta,n}\}$. By definition of mutual information we have,
$I\left(V;\{\mathcal{X}^{(n)},\mathcal{I}^{(n)}\}|Z=1\right)=D_{\text{KL}}\left(\mathbb{P}_{V,\mathcal{X}^{(n)},\mathcal{I}^{(n)}|Z=1}\|\mathbb{P}_{V|Z=1}\mathbb{P}_{\mathcal{X}^{(n)},\mathcal{I}^{(n)}|Z=1}\right)$ (29)
$\stackrel{{\scriptstyle(a)}}{{\leq}}\frac{1}{M^{2}}\sum_{i,j=1}^{M}D_{\text{KL}}\left(\mathbb{P}^{(i)}_{\mathcal{X}^{(n)},\mathcal{I}^{(n)}|Z=1}\|\mathbb{P}^{(j)}_{\mathcal{X}^{(n)},\mathcal{I}^{(n)}|Z=1}\right)$ (30)
where $(a)$ is true, since the KL-divergence is jointly convex. Further we define $\mathbb{P}^{(i)}_{\mathcal{X}^{(n)},\mathcal{I}^{(n)}|Z=1}\triangleq\mathbb{P}_{\mathcal{X}^{(n)},\mathcal{I}^{(n)}|Z=1,V=i}$ to be the conditional distribution of the data given the $i^{\text{th}}$-model is chosen. For any two fixed models $i,j$ we have from Lemma 10 that,
$D_{\text{KL}}\left(\mathbb{P}^{(i)}_{\mathcal{X}^{(n)},\mathcal{I}^{(n)}|Z=1}\|\mathbb{P}^{(j)}_{\mathcal{X}^{(n)},\mathcal{I}^{(n)}|Z=1}\right)\leq\frac{K(\frac{n}{p})\lambda^{2}\log p}{\delta},\quad\text{where K is some universal constant}.$ (31)
Consequently the conditional mutual information is bounded as,
$I(V;\{\mathcal{X}^{(n)},\mathcal{I}^{(n)}\}|Z=1)\leq\frac{K(\frac{n}{p})\lambda^{2}\log p}{\delta}$ (32)
Now substituting for $T=\frac{n}{p}$ and $M=\frac{pd}{4}$, and requiring the maximum probability of error $p_{\text{err}}(\Psi)\leq 1/2$, we get the following lower bound on the observation time,
$T\geq\frac{\log\left(\frac{pd}{4}\right)\delta\left[\frac{2(1-\delta)-1}{2(1-\delta)}\right]}{K\lambda^{2}\log p}$ (33)
Appendix B Proofs of Lemmas
B.1 Proof of Lemma 2 & Lemma 3
As discussed in Section 2.2, the update process can be seen as associating an independent Poisson clock with rate 1 with each random variable, and updating the corresponding random variable at the arrival time accordingly. This view is helpful to calculate the probabilities associated with events $A_{ij}^{k}$ and $D_{ij}^{k}$ for any $i,j\in[p]$ and $k\in[k_{\rm max}]$. We restate and prove both Lemma 2 and 3 in this section.
Lemma 2 (Probability of event $A_{ij}^{k}$).
For $k\in[k_{\rm max}]$, $i,j\in[p]$ we have
$\mathbb{P}\left[A_{ij}^{k}\right]=\left[(1-e^{-\tau/3})e^{-\tau/3}\right]^{3}\triangleq q.$
Proof.
Recall that for $i,j\in[p]$ and any $k\in[k_{\rm max}]$, we define
[…]
$\bigcap\bigg{\{}\exists t\in W_{3}:I^{(t)}=i,\forall t\in W_{3}:I^{(t)}\neq j\bigg{\}}$
where $W_{1}=[(k-1)\tau,(k-2/3)\tau),W_{2}=[(k-2/3)\tau,(k-1/3)\tau),\text{ and }W_{3}=[(k-1/3)\tau,k\tau)$.
The $k-$th $\tau$ length interval being divided into thirds gives us the probability observing at least one update of node $i$ and no updates on node $j$ in the first third as
$\mathbb{P}\left[\exists t\in W_{1}:I^{(t)}=i,\forall t\in W_{1}:I^{(t)}\neq j\right]$ $=\mathbb{P}\left[\exists t\in W_{1}:I^{(t)}=i\right]\cdot\mathbb{P}\left[\forall t\in W_{1}:I^{(t)}\neq j\right]$
$=\mathbb{P}\left[N>0\right]\cdot\mathbb{P}\left[N=0\right]\text{ where }N\sim\operatorname{Pois}(\tau/3)$
$=(1-e^{-\tau/3})e^{-\tau/3}.$
Recall that $I_{t}$ represents the identity of the last updated node on or before time $t$ That is, if $I_{t}=i$ then there exists an update times $S_{n}\leq t<S_{n+1}$ such that $I^{(n)}=i$. As the number of update times (per node) occurring in any interval $\tau$ has Poisson distribution with mean $\tau$ we get the second equality. We get the required result by using independence of the associated Poisson clocks for $W_{2}$ and $W_{3}$ sections of the considered $k$-th interval. ∎
Lemma 3.
The probability of event $D_{ij}^{k}$ for any $k\in[k_{\rm max}]$ we have $\mathbb{P}\left[D_{ij}^{k}\right]\geq e^{-\tau d}.$
Proof.
The proof simply follows from the Poisson distribution and the definition of the event $D_{ij}^{k}$. We have,
$\mathbb{P}\left[D_{ij}^{k}\right]=\mathbb{P}\left[N=0\right]^{\left\lvert N(i)\setminus\{j\}\right\rvert}\geq e^{-\tau d}$
where $N\sim\operatorname{Pois}(\tau)$. ∎
B.2 Proof of Lemma 4
We state and prove Lemma 4 in this section. We use the fact that if the maximum over some $N_{\rm max}>0$ updates of $\epsilon_{I^{(n)}}^{(n)},n\leq N_{\rm max}$ is bounded, then the maximum of the data samples $\mathbf{Y}^{(n)}$ are also bounded by the same quantity up to a factor of $1/(1-d\beta_{\rm max})$. Assumption A3 plays an important role in the proof.
Lemma 4 (Event C).
For any $\delta>0$ and sufficiently large $p$ there exists constants $C_{1},C_{1}^{\prime},C_{1}^{\prime\prime}>0$ such that
$\mathbb{P}\left[\max_{i\in[p],t<C_{1}^{\prime}p^{C_{1}^{\prime\prime}}}\left|Y_{i}^{(t)}\right|\leq C_{1}\sigma_{\rm max}\sqrt{\log\left(\frac{p}{\delta}\right)}\right]\geq 1-\delta/2$
Proof.
We first establish a few claims to prove this lemma.
Claim 1.
Let $N_{\rm max}=T_{\rm max}p=C_{1}^{\prime}p^{C_{1}^{\prime\prime}+1}$ and let $Z=\max_{n\leq N_{\rm max}}\left\lvert\epsilon_{I^{(n)}}^{(n)}\right\rvert$. Then there exists constants $\tilde{C_{1}}$ such that
$\mathbb{P}\left[Z\leq\varepsilon_{\rm max}\right]\geq 1-\delta/2.$
where $\varepsilon_{\rm max}=\tilde{C}_{1}\sigma_{\rm max}\sqrt{\log\left(\frac{p}{\delta}\right)}$.
Proof.
First we bound MGF of $Z$ using the MGF of a folded normal distribution, which is given by $M(t)=2e^{\frac{\sigma^{2}t^{2}}{2}}\Phi(\sigma t)$, where $\Phi(\cdot)$ is the CDF of the standard normal distribution. Therefore,
$\mathbb{E}\left[\exp(tZ)\right]$ $=\mathbb{E}\left[\exp\left(t\max_{n\leq N_{\rm max}}\left\lvert\epsilon_{I^{(n)}}^{(n)}\right\rvert\right)\right]$
=$\stackrel{{\scriptstyle(a)}}{{=}}\mathbb{E}\left[\max_{n\leq N_{\rm max}}\exp\left(t\left\lvert\epsilon_{I^{(n)}}^{(n)}\right\rvert\right)\right]$
<$\stackrel{{\scriptstyle(b)}}{{<}}\sum_{i=1}^{N_{\rm max}}\mathbb{E}\left[\exp\left(t\left\lvert\epsilon_{I^{(n)}}^{(n)}\right\rvert\right)\right]$
$\stackrel{{\scriptstyle(c)}}{{\leq}}N_{\rm max}\left(2e^{\frac{t^{2}\sigma_{\rm max}^{2}}{2}}\Phi(\sigma_{\rm max}t)\right)$
$\stackrel{{\scriptstyle(d)}}{{\leq}}N_{\rm max}\left(2e^{\frac{t^{2}\sigma_{\rm max}^{2}}{2}}\right).$
Equality (a) follows from the exponential being monotonically increasing. Inequality (b) follows from the fact $\max_{i}p_{i}<\sum_{i}p_{i}$ when $p_{i>0}$. Inequality (c) follows from using the MGF of the folded normal distribution and the fact that variance of $\epsilon_{I^{(n)}}^{(n)}$ is atmost $\sigma_{\rm max}^{2}$. While (d) follows from $\Phi(x)\leq 1,x\in\mathbb{R}$.
Next using Chernoff’s technique we get $\Pr(Z\geq\varepsilon)\leq 2N_{\rm max}e^{-\frac{\varepsilon^{2}}{2\sigma_{\rm max}^{2}}}$. Setting this bound to $\delta/2$ implies $\varepsilon>\sigma_{\rm max}\sqrt{2\log\left(\frac{4C_{1}^{\prime}p^{C_{1}^{\prime\prime}+1}}{\delta}\right)}$. Therefore we have,
$\Pr\left\{Z>\sigma_{\rm max}\sqrt{2\log\left(\frac{4C_{1}^{\prime}p^{C_{1}^{\prime\prime}+1}}{\delta}\right)}\right\}\leq\delta/2.$
Let $X^{(n)}_{\rm max}\triangleq\max_{i\in[p]}|X_{i}^{(n)}|$. Now for sufficiently large $p$ there exists constant $\tilde{C}_{1}$ such that
$\max\left\{\sigma_{\rm max}\sqrt{2\log\left(\frac{4C_{1}^{\prime}p^{C_{1}^{\prime\prime}+1}}{\delta}\right)},X^{(0)}_{\rm max}\right\}<\tilde{C}_{1}\sigma_{\rm max}\sqrt{\log\left(\frac{p}{\delta}\right)}.$ (34)
Let $\varepsilon_{\rm max}\triangleq\tilde{C}_{1}\sigma_{\rm max}\sqrt{\log\left(\frac{p}{\delta}\right)}$. Therefore we have,
$\mathbb{P}\left[Z\leq\varepsilon_{\rm max}\right]\geq 1-\delta/2.$
Now we relate this event $Z\leq\varepsilon_{\rm max}$ to the event in the statement of this Lemma using the following claim.
Claim 2.
For $n\leq N_{\rm max},\varepsilon_{\rm max}>0$ the following relation between the events holds:
$\left\{\max_{n\leq N_{\rm max}}\left\lvert\epsilon_{I^{(n)}}^{(n)}\right\rvert<\varepsilon_{\rm max}\right\}\subseteq\left\{\max_{n\leq N_{\rm max}}|X_{I^{(n)}}^{(n)}|\leq\frac{\varepsilon_{\rm max}}{1-d\beta_{\rm max}}\right\}.$ (35)
Proof.
Recall that $\varepsilon_{\rm max}$ chosen in (34) is such that $X^{(0)}_{\rm max}\leq\varepsilon_{\rm max}$, where $X^{(0)}_{\rm max}\geq\left\lvert X_{i}^{(0)}\right\rvert$ for all $i\in[p]$. Suppose for the next round node $i$ was updated, that is $I^{(1)}=i$. Therefore, for node $i$ we have
$\left\lvert X^{(1)}_{i}\right\rvert\leq\sum_{j\in N(i)}\beta_{ij}\left\lvert X_{j}^{(0)}\right\rvert+\left\lvert\epsilon_{i}^{(1)}\right\rvert\leq d\beta_{\rm max}X^{(0)}_{\rm max}+\varepsilon_{\rm max}\leq\varepsilon_{\rm max}(1+d\beta_{\rm max})$
while all other nodes remain unchanged. Then, it follows that
$X_{\rm max}^{(1)}=\max\left\{X_{\rm max}^{(0)},\left\lvert X_{i}^{(1)}\right\rvert\right\}\leq\varepsilon_{\rm max}(1+d\beta_{\rm max}).$
Continuing this way for $n>1$, and using the fact that $d\beta_{\rm max}<1$ from Assumption A3 we get,
$X^{(n)}_{\rm max}$ $<\varepsilon_{\rm max}\left(1+d\beta_{\rm max}+(d\beta_{\rm max})^{2}+\dots+(d\beta_{\rm max})^{n}\right)\leq\frac{\varepsilon_{\rm max}}{1-d\beta_{\rm max}}.$
Now we just combine the two claims and the result follows as
$\mathbb{P}\left[\max_{n\leq N_{\rm max}}|X_{I^{(n)}}^{(n)}|\leq\frac{\varepsilon_{\rm max}}{1-d\beta_{\rm max}}\right]\geq\mathbb{P}\left[Z\leq\varepsilon_{\rm max}\right]\geq 1-\delta/2.$
where $C_{1}=\tilde{C}_{1}/(1-d\beta_{\rm max})$. ∎
B.3 Proof of Lemma 5
In this section, we will state Lemma 5 in the context of definitions from Section 5 and prove it.
Lemma 5 (Conditional probability of event $B_{j}^{k}$).
Under the conditional measure $\mathbb{P}_{C}$ with $\delta>0$, $k\in[k_{\rm max}]$, $i,j\in[p]$ and any vector $\mathbf{\bar{y}}\in\mathbb{R}^{p}$ with $|\mathbf{\bar{y}}|\leq y_{\rm max}$, we have
$\mathbb{P}_{C}\left[B_{j}^{k}\middle|A_{ij}^{k},D_{ij}^{k},\mathbf{Y}^{(k-1)\tau-}=\mathbf{\bar{y}}\right]>0.78-\delta/2\triangleq b.$
Proof.
For convenience, we will drop the subscripts and superscripts on the events and we write the conditional probability as $\mathbb{P}_{y}\left[\cdot\right]=\mathbb{P}\left[\cdot|\mathbf{Y}^{(k-1)\tau-}=\mathbf{\bar{y}}\right]$. The conditional probability of the event $B$ can be written as,
$\mathbb{P}_{C}\left[B|A,D,\mathbf{Y}^{(k-1)\tau-}=\mathbf{\bar{y}}\right]$ $=\frac{\mathbb{P}_{y}\left[B\cap C_{\delta}|A,D\right]}{\mathbb{P}_{y}\left[C_{\delta}|A,D\right]}$
$\stackrel{{\scriptstyle(a)}}{{\geq}}\mathbb{P}_{y}\left[B\cap C_{\delta}|A,D\right]$
=$\stackrel{{\scriptstyle(b)}}{{=}}\mathbb{P}_{y}\left[B|A,D\right]+\mathbb{P}_{y}\left[C_{\delta}|A,D\right]-\mathbb{P}_{y}\left[B\cup C_{\delta}|A,D\right]$
$\stackrel{{\scriptstyle(c)}}{{\geq}}\mathbb{P}_{y}\left[B|A,D\right]+\mathbb{P}_{y}\left[C_{\delta}|A,D\right]-1$
(a) follows by bounding $\mathbb{P}_{y}\left[C_{\delta}|A,D\right]\leq 1$ while (b) follows similarly using $\mathbb{P}_{y}\left[B\cup C_{\delta}|A,D\right]\leq 1$. To lower bound the term $\mathbb{P}_{y}\left[C_{\delta}|A,D\right]$ we use the following claim.
Claim 3.
Under the above conditions on $k,\delta$ and pair $i,j$ we have,
$\mathbb{P}_{y}\left[C_{\delta}|A,D\right]\geq\mathbb{P}\left[C_{\delta}\right]$
Proof.
First note that
$\mathbb{P}\left[C_{\delta}|A,D\right]=\mathbb{P}$ $\left[C_{\delta}\middle|A,D,|\mathbf{Y}^{(k-1)\tau-}|\leq y_{\rm max}\right]\mathbb{P}\left[|\mathbf{Y}^{(k-1)\tau-}|\leq y_{\rm max}\right]$
$+\mathbb{P}\left[C_{\delta}\middle|A,D,|\mathbf{Y}^{(k-1)\tau-}|>y_{\rm max}\right]\mathbb{P}\left[|\mathbf{Y}^{(k-1)\tau-}|>y_{\rm max}\right].$
But by definition of event $C$, we have $\mathbb{P}\left[C_{\delta}\middle|A,D,|\mathbf{Y}^{(k-1)\tau-}|>y_{\rm max}\right]=0$. Therefore,
$\mathbb{P}\left[C_{\delta}|A,D\right]\leq\mathbb{P}\left[C_{\delta}\middle|A,D,|\mathbf{Y}^{(k-1)\tau-}|\leq y_{\rm max}\right]=\mathbb{P}_{y}\left[C_{\delta}\middle|A,D\right].$
Now an implication of Claim 2 is
$C_{\delta}^{c}=\left\{\exists n\leq N_{max},i\in[p]:|X_{i}^{(n)}|>\frac{\varepsilon_{\rm max}}{1-d\beta_{\rm max}}\right\}\subseteq\left\{\exists n\leq N_{max},i\in[p]:\left\lvert\epsilon_{i}^{(n)}\right\rvert>\varepsilon_{\rm max}\right\}.$
This means that, under the conditional measure with respect to the event $A\cap D$ we have,
$\mathbb{P}\left[C_{\delta}^{c}\middle|A,D\right]\leq\mathbb{P}\left[\exists n\leq N_{max},i\in[p]:|\epsilon_{i}^{(n)}|>\varepsilon_{\rm max}\middle|A,D\right].$ (36)
Next we claim that the event $\left\{\exists n\leq N_{max},i\in[p]:|\epsilon_{i}^{(n)}|>\varepsilon_{\rm max}\right\}$ is independent of events $A,D$ that is,
$\mathbb{P}\left[\exists n\leq N_{max},i\in[p]:|\epsilon_{i}^{(n)}|>\varepsilon_{\rm max}\middle|A,D\right]=\mathbb{P}\left[\exists n\leq N_{max},i\in[p]:|\epsilon_{i}^{(n)}|>\varepsilon_{\rm max}\right]$ (37)
To see this, consider the following. Let $\{\epsilon^{(n)}\}_{n=1}^{N_{\rm max}}$ be the Gaussian additions during updates in each round $n\in[N_{\rm max}]$. On the other hand, let the sequence of node identities of updated nodes be $\{I^{(n)}\}_{n=1}^{N_{\rm max}}$.
The update process can be seen as assigning the value of $\epsilon^{(n)}$ to the node $I^{(n)}$ during the update in round $n$. That is, the value of the Gaussian additions can be drawn independently of the identities of the nodes to be updated. Since events $A_{ij}^{k}$ and $D_{ij}^{k}$ depend on updates from the set of nodes $\{i,j\}$ and $N(i)\setminus\{j\}$ respectively, they depend only on the identities of the nodes. Therefore the claim follows. Putting together (37) and (36) we get
$\mathbb{P}\left[C_{\delta}^{c}\middle|A,D\right]\leq\mathbb{P}\left[\exists n\leq N_{max},i\in[p]:|\epsilon_{i}^{(n)}|>\varepsilon_{\rm max}\right]<\delta/2.$
Therefore from the above claim it follows that
$\mathbb{P}_{C}\left[B|A,D,\mathbf{Y}^{(k-1)\tau-}\right]\geq\mathbb{P}_{y}\left[B|A,D\right]+\mathbb{P}_{y}\left[C_{\delta}|A,D\right]-1\geq\mathbb{P}_{y}\left[B|A,D\right]-\delta/2.$
Now we, compute a lower bound on $\mathbb{P}_{y}\left[B|A,D\right]$. Recall that event $B_{j}^{k}$ occurs when
$\left\lvert\Delta Y_{j}^{k}\right\rvert$ $=\left|{Y_{j}^{(k-1/3)\tau}-Y_{j}^{(k-1)\tau}}\right|\geq\sigma_{\rm min}$. Given $\mathbf{Y}^{(k-1)\tau-}=\mathbf{\bar{y}}$ the consequent samples are also Gaussians as linear combination of Gaussian random variables is also a Gaussian and given the event $A_{ij}^{k}$ has occurred, we know the node $j$ was updated at least once in the second third of the interval so the probability of $\left\lvert\Delta Y_{j}^{k}\right\rvert=0$ is zero. Therefore, the random variable $\Delta Y_{j}^{k}$ is Gaussian with some mean $\mu$ and variance $\sigma^{2}\geq 2\sigma_{\rm min}^{2}$. Now, suppose $\mu>0$, then we have
$\mathbb{P}_{y}\left[\left\lvert\Delta Y_{j}^{k}\right\rvert\geq\sigma_{\rm min}\middle|A,D\right]$ $\stackrel{{\scriptstyle(a)}}{{\geq}}\mathbb{P}_{y}\left[\Delta Y_{j}^{k}\geq\sigma_{\rm min}\middle|A,D\right]$
=$\stackrel{{\scriptstyle(b)}}{{=}}\mathbb{P}_{y}\left[Z\geq\frac{\sigma_{\rm min}-\mu}{\sigma}\right]$
$\stackrel{{\scriptstyle(c)}}{{\geq}}\mathbb{P}_{y}\left[Z\geq\frac{\sigma_{\rm min}}{\sqrt{2}\sigma_{\rm min}}\right]$
=$\stackrel{{\scriptstyle(d)}}{{=}}0.7823>0.78$
Equation (b) follows from writing $\Delta Y_{j}^{k}=\sigma Z+\mu$ where $Z\sim\mathcal{N}(0,1)$. Inequality (c) follows from using $\mu>0$ and $\sigma>\sqrt{2}\sigma_{\rm min}$.
B.4 Proof of Lemma 6
In this section we state and prove Lemma 6.
Lemma 6.
For $k\in[k_{\rm max}]$ and $i,j\in[p]$ the $k$-th term of the test statistic is bounded almost surely with respect to the conditional measure $\mathbb{P}_{C}$. That is,
[…]
Proof.
Under the conditional measure $\mathbb{P}_{C}$ we show that the complement of the above event occurs with probability 0. That is,
$\mathbb{P}_{C}\left[\left\lvert T_{ij}^{k}\right\rvert>\frac{2C_{1}\sigma_{\rm max}\sqrt{\log\left(\frac{p}{\delta}\right)}}{\sigma_{\rm min}}\right]$ =$\stackrel{{\scriptstyle(a)}}{{=}}\mathbb{P}_{C}\left[\left\lvert\frac{\Delta Y_{i}^{k}}{\Delta Y_{j}^{k}}\right\rvert>\frac{2C_{1}\sigma_{\rm max}\sqrt{\log\left(\frac{p}{\delta}\right)}}{\sigma_{\rm min}}\middle|AB\right]\mathbb{P}_{C}\left[AB\right]$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\mathbb{P}_{C}\left[\left\lvert\frac{\Delta Y_{i}^{k}}{\sigma_{\rm min}}\right\rvert>\frac{2C_{1}\sigma_{\rm max}\sqrt{\log\left(\frac{p}{\delta}\right)}}{\sigma_{\rm min}}\middle|AB\right]$
$\stackrel{{\scriptstyle(c)}}{{\leq}}\mathbb{P}_{C}\left[\left\lvert\Delta Y_{i}^{k}\right\rvert>2C_{1}\sigma_{\rm max}\sqrt{\log\left(\frac{p}{\delta}\right)}\middle|AB\right]=0$
Equation $(a)$ follows from conditioning on events $A_{ij}^{k}$ and $B_{j}^{k}$ and using the definition of the test statistic. Inequality $(b)$ follows from the fact that given event $B_{j}^{k}$ we have $\left\lvert\Delta Y_{j}^{k}\right\rvert>\sigma_{\rm min}$.
B.5 Proof of Lemma 7 (Separation)
We state Lemma7 and prove it.
Lemma 7 (Separation).
Let $\{\mathcal{F}_{k}\}_{k\geq 1}=\left(\sigma(\{\mathbf{Y}^{(t)}\}_{t=0}^{k\tau})\right)_{k\geq 1}$ be a filtration. For any $k$, if a pair $\{i,j\}\notin E$, we have
$\left\lvert\mathbb{E}_{C}\left[T_{ij}^{k}\middle|\mathcal{F}_{k-1}\right]\right\rvert\leq\frac{2y_{\rm max}}{\sigma_{\rm min}}q\tau d$ (38)
and if $\{i,j\}\in E$,
$\operatorname{sign}(\beta_{ij})\mathbb{E}_{C}\left[T_{ij}^{k}\middle|\mathcal{F}_{k-1}\right]\geq\operatorname{sign}(\beta_{ij})\beta_{\rm min}(1-\tau d)bq-\frac{2y_{\rm max}}{\sigma_{\rm min}}q\tau d.$ (39)
Furthermore, if
$\tau=d^{-1}\left[\frac{6y_{\rm max}}{b\sigma_{\rm min}\beta_{\rm min}}+1\right]^{-1}$
then
$\beta_{\rm min}(1-\tau d)bq=3\left(\frac{2y_{\rm max}}{\sigma_{\rm min}}q\tau d\right).$
Proof.
For convenience, we will fix a pair $i,j\in[p]$ and $k\in[k_{\rm max}]$, and we will write the events without the subscripts or superscripts. Additionally, we write the conditional expectation of the $k^{th}$ test statistic as $\mathbb{E}_{C,\mathcal{F}}\left[T_{ij}\right]=\mathbb{E}_{C}\left[T_{ij}^{k}|\mathcal{F}_{k-1}\right]$. We begin by noting that conditioning on the events $A,B$ and $D$ gives us
$\mathbb{E}_{C,\mathcal{F}}\left[T_{ij}\right]$ $=\mathbb{E}_{C,\mathcal{F}}\left[T_{ij}|ABD\right]\mathbb{P}_{C,\mathcal{F}}\left[ABD\right]+\mathbb{E}_{C,\mathcal{F}}\left[T_{ij}|ABD^{c}\right]\mathbb{P}_{C,\mathcal{F}}\left[ABD^{c}\right].$
This follows from the fact that the test statistic is zero when then events $A_{ij}^{k}$ or $B_{j}^{k}$ do not occur. Additionally, when the event $D$ occurs we have,
$\mathbb{E}_{C,\mathcal{F}}\left[T_{ij}|ABD\right]=\mathbb{E}_{C,\mathcal{F}}\left[\frac{\Delta Y_{i}}{\Delta Y_{j}}|ABD\right]=\mathbb{E}_{C,\mathcal{F}}\left[\beta_{ij}+\frac{\Delta\epsilon_{i}}{\Delta Y_{j}}|ABD\right]=\beta_{ij}.$
That is, the conditional expectation of the test statistic is exactly $\beta_{ij}$. Now consider the case $\{i,j\}\notin E$, then $\beta_{ij}=0$ and we have
$\mathbb{E}_{C,\mathcal{F}}\left[T_{ij}\right]=\mathbb{E}_{C,\mathcal{F}}\left[T_{ij}\middle|ABD^{c}\right]\mathbb{P}_{C,\mathcal{F}}\left[A\right]\mathbb{P}_{C,\mathcal{F}}\left[D^{c}\right]\mathbb{P}_{C,\mathcal{F}}\left[B|A,D^{c}\right].$
The probabilities related to events $A$ and $D^{c}$ factorize as above due to independent uniform selection of nodes during updates. However, these events may not be independent of event $B$, as any updates in the neighborhood of $j$ can affect the distribution (mean) of $X_{j}$. However, in this case we only need an upper bound on the probability of the event $B$ and we use a trivial upper of 1. Therefore we have,
$\left\lvert\mathbb{E}_{C,\mathcal{F}}\left[T_{ij}\right]\right\rvert$ $\stackrel{{\scriptstyle(a)}}{{\leq}}\left\lvert\mathbb{E}_{C,\mathcal{F}}\left[T_{ij}\middle|ABD^{c}\right]\right\rvert q(1-e^{-\tau d})$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\frac{2y_{\rm max}}{\sigma_{\rm min}}q\tau d$
where $(a)$ follows from the fact that $\mathbb{P}_{C,\mathcal{F}}\left[A\right]=\mathbb{P}\left[A\right]$ due to the independent selection of nodes and by definition equal to $q$. Inequality $(b)$ follows from Lemma 6 and Lemma 3 and using $1-e^{-x}<x$. Next, for the case $\{i,j\}\in E$, we have,
$\mathbb{E}_{C,\mathcal{F}}\left[T_{ij}\right]=\beta_{ij}\mathbb{P}_{C,\mathcal{F}}\left[ABD\right]+\mathbb{E}_{C,\mathcal{F}}\left[T_{ij}|ABD^{c}\right]\mathbb{P}_{C,\mathcal{F}}\left[ABD^{c}\right].$
Multiplying $\operatorname{sign}(\beta_{ij})$ both sides and noticing that $\operatorname{sign}(\beta_{ij})\mathbb{E}_{C,\mathcal{F}}\left[T_{ij}\middle|ABD^{c}\right]\leq\left\lvert\mathbb{E}_{C,\mathcal{F}}\left[T_{ij}\middle|ABD^{c}\right]\right\rvert$ we get,
$\operatorname{sign}(\beta_{ij})\mathbb{E}_{C,\mathcal{F}}\left[T_{ij}\right]$ $\geq\operatorname{sign}(\beta_{ij})\beta_{ij}q(e^{-\tau d})b-\left\lvert\mathbb{E}_{C,\mathcal{F}}\left[T_{ij}\middle|ABD^{c}\right]\right\rvert$
$\geq\operatorname{sign}(\beta_{ij})\beta_{\rm min}q(1-\tau d)b-\frac{2y_{\rm max}}{\sigma_{\rm min}}q\tau d.$
This proves the separation in cases where the edge between nodes $i,j$ exists and otherwise. Now for the chosen $\tau=d^{-1}\left[\frac{6y_{\rm max}}{b\sigma_{\rm min}\beta_{\rm min}}+1\right]^{-1}$ we have,
11 11 $=\tau d\left(\frac{6y_{\rm max}}{b\sigma_{\rm min}\beta_{\rm min}}+1\right)$
$1-\tau d$ $=\frac{6y_{\rm max}}{b\sigma_{\rm min}\beta_{\rm min}}\tau d$
$\beta_{\rm min}b(1-\tau d)$ $=3\left[\frac{2y_{\rm max}}{\sigma_{\rm min}}\tau d\right].$
Multiplying $q$ on both sides we have the required result. ∎
B.6 Bernstein-style Concentration Inequality
In the proof of Theorem 1, we use the following Bernstein-style concentration inequality. We state Lemma 8 but we do not prove it here. One can find it here as an implication of [Chung and Lu, 2006, Theorem 27]
Lemma 8 (Bernstein style concentration for martingales).
Let $Z_{1},Z_{2},\dots,Z_{k}$ be a submartingale adapted to filtration $\{\mathcal{F}_{k}\}_{k\geq 1}$ such that $\left\lvert Z_{k}-Z_{k-1}\right\rvert\leq c,c\geq 0$ almost surely and $\operatorname{var}(Z_{k}|\mathcal{F}_{k-1})\leq s^{2},s\geq 0$, then for $t>0$,
$\mathbb{P}\left[Z_{n}<Z_{0}-t\right]\leq\exp\left(-\frac{t^{2}}{2ns^{2}+ct/3}\right)$
B.7 Proof of Lemma 9
We now state and prove Lemma 9 in this section. Recall that $y_{\rm max}$ is bound on data samples and $q$ is the probability of the event $A_{ij}^{k}$ as defined in Section 5.
Lemma 9 (Variance bound).
For an adapted process $\left(Z_{k},\mathcal{F}_{k}\right)_{k\geq 1}$, where $Z_{k}$ is either $\sum_{l=1}^{k}\operatorname{sign}(\beta_{ij})T_{ij}^{l}-k\eta^{\prime}$, $\sum_{l=1}^{k}T_{ij}^{l}-k\eta$ or $\sum_{l=1}^{k}T_{ij}^{l}+k\eta$ and $\{\mathcal{F}_{k}\}_{k\geq 1}=\left(\sigma(\{\mathbf{Y}^{(t)}\}_{t=0}^{k\tau})\right)_{k\geq 1}$, then under the conditional measure $\mathbb{P}_{C}\left[\cdot\right]$ we have
$\operatorname{var}\left(Z_{k}|\mathcal{F}_{k-1}\right)=\frac{4y_{\rm max}^{2}q}{\sigma_{\rm min}}.$
Proof.
Let $D_{k}=Z_{k}-Z_{k-1}$. We have,
$\operatorname{var}\left(Z_{k}|\mathcal{F}_{k-1}\right)$ =$\stackrel{{\scriptstyle(a)}}{{=}}\operatorname{var}\left(Z_{k-1}+D_{k}|\mathcal{F}_{k-1}\right)$
=$\stackrel{{\scriptstyle(b)}}{{=}}\operatorname{var}\left(Z_{k-1}|\mathcal{F}_{k-1}\right)+\operatorname{var}\left(D_{k}|\mathcal{F}_{k-1}\right)+2\operatorname{cov}\left(Z_{k-1},D_{k}|\mathcal{F}_{k-1}\right)$
=$\stackrel{{\scriptstyle(c)}}{{=}}\operatorname{var}\left(D_{k}|\mathcal{F}_{k-1}\right)+2\left[\mathbb{E}_{\mathcal{F}}\left[Z_{k-1},D_{k}\right]-Z_{k-1}\mathbb{E}_{\mathcal{F}}\left[D_{k}\right]\right]$
=$\stackrel{{\scriptstyle(d)}}{{=}}\operatorname{var}\left(D_{k}|\mathcal{F}_{k-1}\right)$
Note that for any $Z_{k}$, the conditional variance of its corresponding difference sequence is
$\operatorname{var}\left(D_{k}|\mathcal{F}_{k-1}\right)$ $=\operatorname{var}\left(T_{ij}^{k}|\mathcal{F}_{k-1}\right)$
$\leq\mathbb{E}_{\mathcal{F}}\left[\left(T_{ij}^{k}\right)^{2}\right]$
$=\mathbb{E}_{\mathcal{F}}\left[\frac{\Delta Y_{i}^{2}}{\Delta Y_{j}^{2}}\middle|AB\right]\mathbb{P}_{\mathcal{F}}\left[AB\right]$
$\leq\frac{4y_{\rm max}^{2}}{\sigma_{\rm min}^{2}}q.$
B.8 Proof of Lemma 10
Note: For convenience of writing in Section 7.2 we excluded the updates times $S_{n}$ from the observations because in the KL-divergence in (40), the joint probability distribution factorizes into $\mathbb{P}_{S_{n}|Z=1}\left[\cdot\right]$ and $\mathbb{P}_{\mathcal{X}^{n},\mathcal{I}^{n}|Z=1}\left[\cdot\right]$ as the Poisson clocks are independent of the uniform selection of the nodes $I^{(n)}$ and the value of the update $X^{(n)}$ at any round $n$. Due to this the variable $S_{n}$ gets integrated out to 1 in the KL-divergence.
Lemma 10.
(Bound on the conditional KL-divergence) Given an observation sequence $\{\mathcal{X}^{(n)},\mathcal{I}^{(n)}\}$ and any two models $\{i,j\}\in[M]$, KL-divergence between the conditional data distributions is upper bounded as,
$D_{\text{KL}}\left(\mathbb{P}^{(i)}_{\mathcal{X}^{(n)},\mathcal{I}^{(n)}|Z=1}\middle\|\mathbb{P}^{(j)}_{\mathcal{X}^{(n)},\mathcal{I}^{(n)}|Z=1}\right)\leq\left(\frac{n}{p}\right)\frac{K\lambda^{2}\log p}{\delta}$ (40)
where $\mathbb{P}^{(i)}_{\cdot}=\mathbb{P}_{\cdot|V=i}$ for any $i\in[M]$.
Proof.
Let $\{\mathcal{X}^{(n)},\mathcal{I}^{(n)}\}=\{\mathbf{X}^{(1)},\ldots,\mathbf{X}^{(n)},I^{(2)},\ldots,I^{(n)}\}$ be a given observation sequence, where $X^{(1)}$ is a Gaussian random vector drawn according to some stationary distribution $\mathcal{N}(0,\Theta^{-1})$ and $\{\mathbf{X}^{(2)},\dots,\mathbf{X}^{(n)}\}$ are n vector samples that evolve according to some Glauber process. Recall that the node update indices are given by $\{I^{(2)},\ldots,I^{(n)}\}$, with each node picked uniformly at random. The conditional data distribution is then given by,
$\mathbb{P}^{(i)}_{\mathcal{X}^{(n)},\mathcal{I}^{(n)}|Z=1}$ =$\stackrel{{\scriptstyle(a)}}{{=}}\mathbb{P}^{(i)}\left(\mathcal{I}^{(n)}|Z=1\right)\mathbb{P}^{(i)}\left(\mathcal{X}^{(n)}|Z=1,\mathcal{I}^{(n)}\right)$ (41)
=$\stackrel{{\scriptstyle(b)}}{{=}}\mathbb{P}^{(i)}\left(I^{(2)},\ldots,I^{(n)}\right)\mathbb{P}^{(i)}\left(\{\mathbf{X}^{(1)},\dots,\mathbf{X}^{(n)}\}|Z=1,\mathcal{I}^{(n)}\right)$ (42)
=$\stackrel{{\scriptstyle(c)}}{{=}}\left(\frac{1}{p}\right)^{n-1}\mathbb{P}^{(i)}\left(\mathbf{X}^{(1)}\right)\prod_{l=2}^{n}\mathbb{P}^{(i)}\left(\mathbf{X}^{(l)}|Z=1,\mathbf{X}^{(l-1)},I^{(l)}\right).$ (43)
where $(a)$ follows from the chain rule of probability, $(b)$ follows since the node update process is independent of the event $\{Z=1\}$ and $(c)$ is true by chain rule and noting that the node update process is uniformly random. Subsequently, the KL-divergence between the conditional distribution |$\mathbb{P}^{(i)}_{\cdot|Z=1}$ and |$\mathbb{P}^{(j)}_{\cdot|Z=1}$ decomposes as,
$D_{\text{KL}}\left(\mathbb{P}^{(i)}_{\mathcal{X}^{(n)},\mathcal{I}^{(n)}|Z=1}\|\mathbb{P}^{(j)}_{\mathcal{X}^{(n)},\mathcal{I}^{(n)}|Z=1}\right)=C_{1}+\sum_{l=2}^{n}C_{l}$ (44)
where,
$C_{1}=D_{\text{KL}}\left(\mathbb{P}^{(i)}_{X^{(1)}}\|\mathbb{P}^{(j)}_{X^{(1)}}\right)\quad\text{and}\quad C_{l}=D_{\text{KL}}\left(\mathbb{P}^{(i)}_{{X}^{(l)}|Z=1,X^{(l-1)},I^{(l)}}\|\mathbb{P}^{(j)}_{X^{(l)}|Z=1,X^{(l-1)},I^{(l)}}\right)$ (45)
We proceed by separately bounding the terms $C_{1}$ and $C_{l}$. We start with $C_{1}$, note that by definition, $\mathbb{P}^{(i)}_{X}$ is distributed as $\mathcal{N}(0,\Theta^{-1}_{i})$ and similarly, $\mathbb{P}^{(j)}_{X}$ as $\mathcal{N}(0,\Theta^{-1}_{j})$. Furthermore, by the construction of the restricted ensemble, the corresponding graphs $G_{i}$ and $G_{j}$ differ only by two edges, with the edge strength denoted by $\lambda$. This implies that $\Theta_{i}$ is $\lambda$-perturbed from $\Theta_{j}$, and vice versa. The KL-divergence for such models is known and is given by $C_{1}\leq 4\lambda^{2}$, (see proof of Theorem 2 in Wang et al. [2010]). We now proceed to bound $C_{l}$. Recall from the construction of the restricted ensemble that the graphs $G_{i}$ and $G_{j}$ differ exactly at two edges, denoted by $e_{i}$ and $e_{j}$. By this construction, the KL-divergence $C_{l}$ is nonzero only if the node update index $I^{(l)}\in e_{i}\cup e_{j}$ otherwise, $C_{l}=0$. This is because the neighborhood of $X_{I^{(l)}}$ remains unchanged under both models unless $I^{(l)}$ is directly involved in the edges where the graphs differ. Consequently, we have
$C_{l}$ $=C_{l|I^{(l)}\in e_{i}\cup e_{j}}\cdot P(I^{(l)}\in e_{i}\cup e_{j})$ (46)
$=\frac{4}{p}C_{l|I^{(l)}=t},\quad\text{where $u\in e_{i}$}$ (47)
where $C_{l|I^{(l)}=t}=D_{\text{KL}}\left(\mathbb{P}^{(i)}_{{X}^{(l)}|Z=1,X^{(l-1)},I^{(l)}=t}\|\mathbb{P}^{(j)}_{X^{(l)}|Z=1,X^{(l-1)},I^{(l)}=t}\right)$. Now it remains to compute
$C_{l|I^{(l)}=t}$. Notice that the distribution $\mathbb{P}^{(i)}\left(X^{(l)}|X^{(l-1)},I^{(l)}=t\right)$ is just the conditional distribution given by $\mathbb{P}^{(i)}\left(X_{t}|X_{[p]/\{t\}}\right)$, here the superscripts containing the time point $l$ is relaxed for ease of notation. From the Glauber update process, we have:
$X_{t}$ $=\sum_{k\neq t}\beta^{(i)}_{tk}X_{k}+\epsilon^{(i)}_{t},\quad\text{where $\epsilon^{(i)}_{t}\sim\mathcal{N}(0,\sigma^{2(i)}_{t})$}$ (48)
Therefore, $\mathbb{P}^{(i)}\left(X_{t}|X_{[p]/\{t\}}\right)\sim\mathcal{N}(\mu^{(i)}_{t},\sigma^{2(i)}_{t})$, where $\mu^{(i)}_{t}$ and $\sigma^{2(i)}_{t}$ are the conditional mean and variance given by:
$\mu^{(i)}_{t}=\sum_{k\neq t}\beta^{(i)}_{tk}X_{k},\quad\text{and}\quad\sigma^{2(i)}_{t}=\frac{1}{\Theta^{(i)}_{tt}}.$ (49)
Furthermore, recall that conditioning on the event $\{Z=1\}$ implies that $\max_{i\in[n]}|X^{(i)}|<C\sqrt{\frac{\log n}{\delta}}$. Consequently, the resulting distribution $\mathbb{P}^{(i)}_{X_{t}|Z=1,X_{[p]/\{t\}}}$ conforms to a truncated Gaussian distribution, with truncation occurring at the value $C\sqrt{\frac{\log n}{\delta}}$. The quantity $C_{l|I^{(l)}=t}$ captures the KL-divergence between two truncated Gaussian distributions which are truncated at the same point, i.e., $C\sqrt{\frac{\log n}{\delta}}$. To compute this, we use well-known results on the KL-divergence between truncated Gaussians (see Nielsen [2022]). We therefore have,
$C_{l|I^{(l)}=t}=\log\frac{Z^{(j)}_{t}}{Z^{(i)}_{t}}+\frac{1}{2}\left[\log\frac{\sigma^{(j)}_{t}}{\sigma^{(i)}_{t}}-\frac{\text{Var}_{\phi^{(i)}_{t}}}{2\sigma^{2(j)}_{t}}+\frac{\text{Var}_{\phi^{(i)}}+(\mu^{i}_{t}-\mu^{j}_{t})^{2}}{2\sigma^{2(i)}_{t}}\right]$ (50)
where $Z^{(i)}_{t}$ and $Z^{(j)}_{t}$ are the normalization functions of the truncated Gaussian, and $\text{Var}_{\phi^{(i)}_{t}}$ is the variance of the truncated Gaussian, with $\phi^{(i)}_{t}=\mathbb{P}^{(i)}_{X_{t}|Z=1,X_{[p]/\{t\}}}$. The expression in equation (50) simplifies significantly because $\sigma^{2(i)}_{t}=\sigma^{2(j)}_{t}$. By the construction of the restricted ensemble, we have $\sigma^{2(i)}_{t}=1/\Theta^{(i)}_{tt}$, since the diagonals of $\Theta^{(i)}=1+\lambda$ for all $i\in[M]$. Furthermore since $\Theta^{(i)}$ and $\Theta^{(j)}$ differ in only one edge, we have by definition $(\mu^{i}_{t}-\mu^{j}_{t})^{2}\leq K^{\prime}\lambda^{2}\frac{\log p}{\delta}$ for some constant $K^{\prime}$. Putting these together, we get
$C_{l}=\frac{4}{p}C_{l|I^{l}=t}\leq\frac{2K^{\prime}\lambda^{2}\log p(1+\lambda)^{2}}{2\delta}\leq\frac{K^{\prime}\lambda^{2}\log p}{\delta}$ (51)
Finally, we have,
$D_{\text{KL}}\left(\mathbb{P}^{(i)}_{\mathcal{X}^{(n)},\mathcal{I}^{(n)}|Z=1}\|\mathbb{P}^{(j)}_{\mathcal{X}^{(n)},\mathcal{I}^{(n)}|Z=1}\right)$ $=C_{1}+\sum_{l=2}^{n}C_{l}$ (52)
$\leq 4\lambda^{2}+\left(\frac{n-1}{p}\right)\frac{K^{\prime}\lambda^{2}\log p}{\delta}$ (53)
$\leq\left(\frac{n}{p}\right)\frac{K\lambda^{2}\log p}{\delta}$ (54)
Appendix C Convergence Rate of the Single-Site Gibbs Sampler
Consider a $p$-dimensional multivariate normal distribution $\mathcal{N}(0,\Sigma)$, where $\Sigma$ is a positive definite covariance matrix. Suppose we are interested in obtaining $m$ approximately independent samples from this distribution using the single-site Gibbs sampler. The single-site Gibbs sampler updates one coordinate at a time, sampling from the conditional distribution of each variable given the others. The convergence rate of this Markov chain depends on both the condition number of the covariance matrix and the dimension $p$. Let $\lambda_{\max}$ and $\lambda_{\min}$ denote the largest and smallest eigenvalues of $\Sigma$, respectively, and let $\kappa(\Sigma)=\lambda_{\max}/\lambda_{\min}$ be the condition number of $\Sigma$.
The mixing time $t_{\text{mix}}(\epsilon)$, defined as the number of iterations required for the total variation distance to the stationary distribution to be less than $\epsilon$, satisfies the following bound:
$t_{\text{mix}}(\epsilon)\geq C\cdot p\cdot\kappa(\Sigma)\cdot\log\left(\frac{1}{\epsilon}\right),$ (55)
where $C$ is an absolute constant [Roberts and Sahu, 1997]. This follows from simple bounds on the spectral gap $\gamma$ of the Markov transition operator underlying the sampler [Liu et al., 1994; Amit, 1991]. This implies that in high-dimensional settings, the mixing time increases linearly with the dimension $p$ and the condition number $\kappa(\Sigma)$. Therefore, to obtain $m$ approximately independent samples, we need to run the Gibbs sampler for at least $\Omega\left(mp\kappa(\Sigma)\right)$, up to log factors. We know that $m\sim d^{2}\log p$ samples are needed for graphical model selection using a standard method such as the Graphical LASSO Yuan and Lin [2007]; Ravikumar et al. [2011]. On the other hand, our algorithm learns the graph in order $pd^{2}\operatorname{poly}\log(p)$ time, which (upto $\operatorname{poly}\log$ factors) is comparable to the samples required for mixing and applying traditional algorithms.